### v1

# Monotone Neural Barrier Certificates

Abstract
This work presents a neurosymbolic framework for safety verification and control synthesis in high-dimensional monotone dynamical systems, using an algorithm with linear sample complexity—without relying on explicit models or conservative Lipschitz bounds. The approach combines the expressiveness of neural networks with the rigor of symbolic reasoning via barrier certificates, functional analogs of inductive invariants that formally guarantee safety. Prior data-driven methods often treat dynamics as black-box models, relying on dense state-space discretization or Lipschitz overapproximations, leading to exponential sample complexity. In contrast, monotonicity—a pervasive structural property in many real-world systems—provides a symbolic scaffold that simplifies both learning and verification. Exploiting order preservation reduces verification to localized boundary checks, transforming a high-dimensional problem into a tractable, low-dimensional one. Barrier certificates are synthesized using monotone neural networks—architectures with embedded monotonicity constraints—trained via gradient-based optimization guided by barrier conditions. This enables scalable, formally sound verification directly from simulation data, bridging black-box learning and formal guarantees within a unified neurosymbolic framework. Empirical results on three large-scale benchmarks—a freeway traffic model with increasing size, a 50-dimensional urban traffic network, and a 13,000-dimensional power grid—demonstrate the scalability and effectiveness of the approach in real-world, safety-critical systems.
1 Introduction
Discrete-time dynamical systems offer a foundational framework for modeling and controlling a broad spectrum of digitally controlled autonomous systems, including autonomous vehicles, implantable medical devices, biological networks, and industrial process management. As these systems constitute a critical component of our safety infrastructure, high-assurance formal reasoning is essential to ensure their safe operation. A principled and efficient approach to providing such safety guarantees involves synthesizing barrier certificates prajna2006barrier ; ADA-SC-ME-GN-KS-PT:19 —a functional analog of inductive invariants—whose existence serves as a formal proof of safety AP-AR:04 . Although finding such certificates is inherently computationally challenging, recent neuro-symbolic approaches have shown considerable promise through the use of neural barrier certificates chang2019neural ; zhao2020synthesizing ; zhang2024exact ; robey2020learning . While these neural certificates provide substantial flexibility in searching for safety guarantees, we posit that scalability and sample efficiency can be further improved in specialized control applications by identifying and leveraging structural properties of the underlying system. In this work, we focus on a subclass of discrete-time dynamical systems known as monotone systems angeli2003monotone ; HLS:95 , which frequently arise in domains such as queuing networks farina2011positive , population dynamics leenheer2004predator , biological systems DA:13 , and chemical networks EDS:07 . By leveraging their structural characteristics, we synthesize barrier certificates represented as monotone neural networks, which we refer to as monotone neural barrier certificates.
Monotone Systems. Monotone dynamical systems HLS:95 ; angeli2003monotone are characterized by the property that their trajectories preserve a natural partial order induced on their state set. As a paradigmatic example, consider traffic flow dynamics sadraddini2018formal , where increasing the number of vehicles on a road segment does not lead to a decrease in vehicle density downstream under normal flow conditions. That is, if one traffic configuration contains more vehicles than another, the evolution of the system preserves this ordering. Similar examples arise in domains such as biochemical reactions in synthetic biology EDS:07 , frequency regulation in power system, congestion control in transportation networks SC-MA:15 ; GC-EL-KS:15 , population dynamics in ecology, and compartmental models in epidemiology angeli2003monotone . Related concepts have also been studied in other domains, including positive linear systems farina2011positive , well-structured transition systems FINKEL200163 , and well-behaved transition systems Blondin2016WellBT . It is well known that monotone systems exhibit highly ordered transient and asymptotic behaviors HLS:95 , making them particularly well-suited for data-driven methods, as their inherent structure helps mitigate the challenges posed by limited or noisy data. While monotonicity has informed control and stability analysis AR:15 ; SC:19 , its potential for learning-based safety verification remains largely untapped.
Barrier Certificates (A Proof of Safety). In safety verification, the central object of interest is the reachable set, which comprises all states the system can reach from a given set of initial states under a fixed control policy. Safety can be certified by showing that this set does not intersect the set of harmful or unsafe states. However, computing the exact reachable set is generally intractable for nonlinear or high-dimensional systems. A common alternative is to construct an inductive state-invariant, i.e., a set that contains the initial states and is closed under the system’s transitions. Barrier certificates prajna2006barrier offer one such approach by defining a real-valued function over the state space whose level sets form a “barrier” separating safe and unsafe regions. Specifically, the barrier function must: 1) take values less than or equal to a constant over the initial states, 2) and strictly greater than that constant over unsafe states; and 3) decrease along system trajectories outside the unsafe set. A control barrier certificate (CBC) generalizes the idea of barrier certificate by requiring the existence of an action satisfying the conditions for a barrier certificate.
Despite their intuitive appeal and ease of verification, complexity of finding CBCs remains a significant hurdle in their widespread adoption. Traditional methods for synthesizing CBCs rely on optimization-based formulations such as Sum-of-squares (SOS) programming AP-SJ:05 ; kong2013exponential . SOS is a popular approach that is applicable to systems with polynomial dynamics; however, it quickly becomes intractable due to the growth of semidefinite constraints with system dimension and polynomial degree. Linear programming (LP) methods, while more tractable, are typically limited to linear or affine certificates and cannot capture the rich nonlinear boundaries present in many real-world systems SS-XC-EA:13 ; YZ-etal:16 . Recent progress in deep learning has enabled the use of neural networks to be trained as barrier certificates anand2023formally ; lindemann2024learning ; dawson2022safe ; liu2023safe ; zhao2020synthesizing ; Rus2023 ; zhang2024exact . These so-called neural control barrier certificates (NCBCs) represent both the barrier function and controller as neural networks. While expressive, these approaches often require full knowledge of the system dynamics (to verify correctness using constraint satisfaction tools such as Z3 de2008z3 ) or rely on bounding Lipschitz constants (which entails exponential sample complexity), limiting scalability and applicability to black-box systems anand2023formally ; zhang2024exact . Moreover, these approaches treat the system as unstructured, ignoring properties like monotonicity that could drastically reduce the number of required samples and constraints [^1] We refer to Appendix A for a comprehensive review of related works..
Monotone Neural Control-Barrier Certificates. We introduce a machine learning algorithm with linear sample complexity for synthesizing control barrier certificates specifically tailored to (unknown) monotone systems. By exploiting monotonicity, we restrict the satisfaction of safety conditions to critical boundary slices of the state space—namely, the upper boundary of the initial set and the lower boundary of the unsafe set. This reduces the verification burden from the entire state space to a carefully selected subset—without compromising formal guarantees—and enables the use of gradient-based optimization to learn barrier certificates using simulation data alone. Our method avoids the need for explicit models, Lipschitz bounds, or semidefinite relaxations, and yields an algorithm for synthesizing barrier certificates that scales linearly with the system dimension. This makes it uniquely suited for safety verification in large-scale, nonlinear systems where monotonicity arises naturally from physical or structural properties.
Contributions. In summary, this work introduces a learning framework for safety verification and safe control synthesis of unknown monotone dynamical systems using monotone neural barrier certificates. The key contributions of our paper are as follows:
  1. 1.
    Monotone structure of barrier certificates. We investigate the structure of barrier certificates for monotone dynamical systems and establish a theoretical result demonstrating that, under mild conditions on the unsafe set, monotone barrier certificates can effectively characterize the system’s safety. This result shows that restricting to monotone certificates incurs no loss of generality and supports their use to improve sample complexity and computational cost of safety verification.
  2. 2.
    Monotonicity as a catalyst for scalable verification. By leveraging the monotone structure of barrier certificates, we propose a scalable framework that verifies safety using only a finite amount of simulator data—without requiring explicit models or Lipschitz constants—by reducing verification to localized boundary checks.
  3. 3.
    Monotone Neural Barrier Certificates. We propose an algorithm that employs monotone neural networks to learn barrier certificates by leveraging reduced safety conditions on the boundaries of the state set. Unlike existing methods based on sum-of-squares, linear programming, or neural barrier certificates, our algorithm achieves linear sample complexity—rather than exponential—and supports arbitrary partitioning of the state set.
  4. 4.
    Validation on large-scale, safety-critical domains. We evaluate the efficiency of our algorithm on three challenging case studies: (i) a 1,000-dimensional freeway traffic model, (ii) a 50-dimensional urban traffic control network, and (iii) a 13,000-dimensional power grid. These results demonstrate the practicality and scalability of our method for safety-critical real-world applications. Moreover, in the high-dimensional freeway traffic model, we show that our algorithm outperforms state-of-the-art methods for synthesizing barrier certificates.
2 Preliminaries and Mathematical Foundations
Notations.
We denote the set of real, non-negative real, and positive real numbers by $\mathbb{R}$, $\mathbb{R}_{\geq 0}$, and $\mathbb{R}_{>0}$, respectively. For sets $A$ and $B$, we use $A\setminus B$ for set difference and $A\times B$ for Cartesian product, and write $|A|$ for the cardinality of $A$. Given a vector $v=(v_{1},\ldots,v_{n})^{\top}\in\mathbb{R}^{n}$, its Euclidean norm is $\|v\|=\sqrt{\sum_{i=1}^{n}v_{i}^{2}}$. The rectified linear unit is defined by $\mathsf{ReLU}(x):=\max(x,0)$, and for vectors $x,y\in\mathbb{R}^{n}$, the mean squared error is $\mathsf{MSE}(x,y):=\frac{1}{n}\sum_{i=1}^{n}(x_{i}-y_{i})^{2}$. We denote the $i$th component of a vector $x$ by $x_{i}$, and write $x\leq y$ to indicate element-wise inequality. A set $\mathcal{S}\subseteq\mathbb{R}^{n}$ is lower closed if $x\in\mathcal{S}$ and $y\leq x$ imply $y\in\mathcal{S}$; it is upper closed if $x\in\mathcal{S}$ and $y\geq x$ imply $y\in\mathcal{S}$. A hyper-rectangle is denoted by $\underline{x},\overline{x}]:=\{x\in\mathbb{R}^{n}\mid\underline{x}\leq x\leq\overline{x}\}$. Given a set $A\subseteq^{n}$, $x\in^{n}$ is a maximal (resp. minimal) point of $A$, if there does not exists any $y\in A$ such that $x\leq y$ (resp. $y\leq x$).
2.1 Safety of Dynamical Systems
In this section, we provide a formal definition of discrete-time dynamical systems and establish a rigorous notion of safety. We then present two central safety objectives.
Definition 1 (Discrete-time dynamical system).
A discrete-time control system (dtDS) is a tuple $\mathfrak{S}:=(\mathcal{X},\mathcal{X}_{0},U,f)$, where $\mathcal{X}\subseteq\mathbb{R}^{n}$ represents the state set, $\mathcal{X}_{0}\subseteq\mathcal{X}$ is the initial state set, $U\subseteq\mathbb{R}^{m}$ is the set of inputs, and $f:\mathcal{X}\times U\to\mathcal{X}$ is the state transition map.
Given an input sequence $\{u_{t}\}_{t=0}^{\infty}\in U$, the evolution of $\mathfrak{S}$ under this input is described by:
$x_{t+1}=f(x_{t},u_{t}),\qquad\text{for all }t\geq 0.$ (1)
The sequence $\{x_{t}\}_{t=0}^{\infty}$ defined by (1) is called a trajectory of $\mathfrak{S}$ for the input $\{u_{t}\}_{t=0}^{\infty}$. Given a dtDS $\mathfrak{S}=(\mathcal{X},\mathcal{X}_{0},U,f)$, a set of unsafe states $\mathcal{X}_{u}\subseteq\mathcal{X}$, and an input sequence $\{u_{t}\}_{t=0}^{\infty}$, we say that $\mathfrak{S}$ is safe with respect to $\mathcal{X}_{u}$ for this input sequence if all trajectories starting from initial states in $\mathcal{X}_{0}$ remain outside of $\mathcal{X}_{u}$. That is, for every $x_{0}\in\mathcal{X}_{0}$, the corresponding trajectory $\{x_{t}\}_{t=0}^{\infty}$ for the input sequence $\{u_{t}\}_{t=0}^{\infty}$ starting from $x(0)=x_{0}$, satisfies
$\{x_{t}\}_{t=0}^{\infty}\cap\mathcal{X}_{u}=\emptyset.$
In dynamical systems, how inputs are managed fundamentally shapes safety objectives. Hence, two core challenges arise: the safety verification problem and the safe control synthesis problem.
Safety Verification problem.
In safety verification, inputs are viewed as disturbances, and the goal is to guarantee system safety under all possible input sequences $\{u_{t}\}_{t=0}^{\infty}\in U$. To this end, robust barrier certificates have been widely employed as an effective tool for safety verification (prajna, ).
Definition 2 (Robust barrier certificates).
Consider a dtDS $\mathfrak{S}=(\mathcal{X},\mathcal{X}_{0},U,f)$ with a set of unsafe states $\mathcal{X}_{u}\subseteq\mathcal{X}$. A function $B:\mathcal{X}\rightarrow\mathbb{R}$ is called a robust barrier certificate (RBC) for $\mathfrak{S}$ with respect to the unsafe set $\mathcal{X}_{u}\subseteq\mathcal{X}$ if there exists $\gamma,\delta,\eta\in\mathbb{R}_{\geq 0}$ such that $\gamma\leq\delta\leq\eta$, and:
$B(x)\leq\gamma,\qquad\text{ for all }x\in\mathcal{X}_{0},$ (2)
$B(x)>\eta,\qquad\text{ for all }x\in\mathcal{X}_{u},\text{ and }$ (3)
$B(x)\leq\delta\implies\mathrm{sup}_{u\in U}\{B(f(x,u))\}\leq\delta.$ (4)
The following theorem formulates the safety verification problem as a search for RBCs prajna .
Theorem 1 (Safety verification via robust barrier certificates).
Consider a dtDS $\mathfrak{S}=(\mathcal{X},\mathcal{X}_{0},U,f)$ and a set of unsafe states $\mathcal{X}_{u}\subseteq\mathcal{X}$. Suppose that there exists a robust barrier certificate satisfying conditions (2)-(4). For any input sequence $\{u_{t}\}_{t=0}^{\infty}\in U$, the system $\mathfrak{S}$ is safe with respect to $\mathcal{X}_{u}$.
Proof. We refer to Appendix B for the proof.
Safe Control Synthesis problem.
In safe control synthesis, inputs are treated as control policies, aiming to synthesize a controller $\{u_{t}\}_{t=0}^{\infty}$ that ensures system safety. Control barrier certificates have been widely adopted as an effective tool for synthesizing such controllers (ADA-XX-JWG-PT:17, ).
Definition 3 (Control barrier certificates).
Consider a dtDS $\mathfrak{S}=(\mathcal{X},\mathcal{X}_{0},U,f)$ with a set of unsafe states $\mathcal{X}_{u}\subseteq\mathcal{X}$. A function $B:\mathcal{X}\rightarrow\mathbb{R}$ is called a Control Barrier Certificate (CBC) for $\mathfrak{S}$ with respect to the unsafe set $\mathcal{X}_{u}\subseteq\mathcal{X}$ if there exists $\gamma,\delta,\eta\in\mathbb{R}_{\geq 0}$ such that $\gamma\leq\delta\leq\eta$ and $B$ satisfy conditions (2), (3), and the following condition
$B(x)\leq\delta\implies\mathrm{inf}_{u\in U}\{B(f(x,u))\}\leq\delta.$ (5)
The next result formulates the safe control synthesis problem as a search for CBCs ADA-XX-JWG-PT:17 .
Theorem 2 (Control synthesis via control barrier certificates).
Consider a dtDS $\mathfrak{S}=(\mathcal{X},\mathcal{X}_{0},U,f)$ and a set of unsafe states $\mathcal{X}_{u}\subseteq\mathcal{X}$. Suppose that there exists a control barrier certificate satisfying conditions (2), (3), and (5). Then, there exists a map $K:\mathcal{X}\to U$ such that the system $\mathfrak{S}$ with the feedback controller $\{u_{t}=K(x_{t})\}_{t=0}^{\infty}$ is safe with respect to $\mathcal{X}_{u}$.
Proof. We refer to Appendix C for the proof.
2.2 Monotone Dynamical Systems
In this subsection, we introduce a special class of discrete-time dynamical systems known as monotone systems. We begin by formally defining the notion of monotonicity.
Definition 4 (Monotone maps).
Consider the sets $\mathcal{X}{\subseteq}\mathbb{R}^{n}$ and $\mathcal{Y}{\subseteq}\mathbb{R}^{m}$. Then the map $f:\mathcal{X}{\to}\mathcal{Y}$ is monotone if, for every $x,y\in\mathcal{X}$ such that $x\leq y$, we have $f(x)\leq f(y)$.
Using this notion of monotonicity, we define monotone dynamical systems.
Definition 5 (Monotone dynamical systems).
A dtDS $\mathfrak{S}=(\mathcal{X},\mathcal{X}_{0},U,f)$ is monotone if the state transition map $f:\mathcal{X}\times U\to\mathcal{X}$ is monotone.
For a differentiable transition map $f$ and convex sets $\mathcal{X}$ and $U$, the system $\mathfrak{S}$ is monotone if and only if $Df(x,u)\geq 0_{n\times(n+m)}$ for all $(x,u)\in\mathcal{X}\times U$ HLS:95 . Monotone systems preserve the standard partial order along their trajectories HLS:95 ; angeli2003monotone , resulting in highly regular behaviors that facilitate reasoning about dynamics via partial orders.
Problem Formulation.
We consider the safety of a monotone dtDS $\mathfrak{S}=(\mathcal{X},\mathcal{X}_{0},U,f)$ with an unknown transition map $f$. While $f$ is not explicitly known, we assume access to a simulator that returns $f(x,u)$ for any $x\in\mathcal{X}$ and $u\in U$. This paper focuses on the following problem.
Problem 1 (Safety Certificates for Unknown Monotone Systems).
Consider a monotone dtDS $\mathfrak{S}=(\mathcal{X},\mathcal{X}_{0},U,f)$ with an unknown state-transition map $f$, and a set of unsafe states $\mathcal{X}_{u}\subseteq\mathcal{X}$. Our goal is to use a finite set of samples from a simulator of $f$ to:
  1. 1.
    (Safety verification): Verify that $\mathfrak{S}$ is safe with respect to $\mathcal{X}_{u}$ for all input $\{u_{t}\}_{t=0}^{\infty}\in U$.
  2. 2.
    (Safe control synthesis): Design a feedback controller $K:\mathcal{X}\rightarrow U$ such that $\mathfrak{S}$ with the feedback controller $\{u_{t}=K(x_{t})\}_{t=0}^{\infty}$ is safe with respect to $\mathcal{X}_{u}$.
The monotonicity assumption in Problem 1, although made in the absence of an explicit model, is often justified by prior knowledge of the system’s physical properties. Many physical processes naturally exhibit monotonicity, where increases in inputs or states lead to non-decreasing responses. This property is common in domains such as traffic networks SC-MA:15 , power systems park2020uniqueness , biological processes EDS:07 , and supply chains yu2019equilibrium .
3 Monotone Neural Barrier Certificates
We propose a data-driven framework that leverages system monotonicity to learn barrier functions using neural networks. This approach enables the construction of sound safety certificates from limited data, avoiding full system identification or exhaustive sampling. We first analyze the structure of barrier certificates for monotone systems and show that, under certain assumptions on the unsafe set $\mathcal{X}_{u}$, any safe monotone system admits a monotone robust (resp. control) barrier certificate.
3.1 Monotone Barrier Certificates
In this subsection, we analyze the structure of barrier certificates for safety verification and control synthesis in a monotone system $\mathfrak{S}=(\mathcal{X},\mathcal{X}_{0},U,f)$. We show that if the unsafe set $\mathcal{X}_{u}$ is upper closed, both problems can be addressed using monotone barrier certificates.
Theorem 3 (Monotone Barrier Certificates).
Consider a monotone dtDs $\mathfrak{S}:=(\mathcal{X},\mathcal{X}_{0},U,f)$ with an upper-closed set of unsafe states $\mathcal{X}_{u}\subseteq\mathcal{X}$. Then, the following statements hold:
  1. 1.
    (Safety verification). $\mathfrak{S}$ is safe with respect to $\mathcal{X}_{u}$, for every input $\{u_{t}\}_{t=0}^{\infty}\in U$ if and only if there exists a monotone robust barrier certificate satisfying (2),(3), and (4).
  2. 2.
    (Safe control synthesis). there exists a map $K:\mathcal{X}\to U$ such that $\mathfrak{S}$ with the feedback controller $\{u_{t}=K(x_{t})\}_{t=0}^{\infty}$ is safe with respect to $\mathcal{X}_{u}$ if and only if there exists a monotone control barrier certificate satisfying (2),(3), and (5).
Proof. We refer to Appendix D for the proof.
For monotone systems, Theorem 3 reduces the search for robust (control) barrier certificates to monotone barrier functions. As will be shown in Section 3.2, the monotonicity structure of barrier certificates plays a crucial role in our framework, leading to significantly lower sample complexity and computational cost—particularly in high-dimensional systems. This structure also enables restricting learning-based approximators to inherently monotone functions. Monotone neural networks thus provide a natural and expressive choice for approximating such certificates.
Monotone Neural Networks.
We consider a neural network $N$ with $k$ fully connected layers, where each layer $i$ is defined by a weight matrix $W_{i}$ and a bias vector $b_{i}$ of appropriate dimensions, followed by an activation function $\sigma$. The network represents a function $N:\mathbb{R}^{n_{0}}\to\mathbb{R}^{n_{k}}$. Given an input $u\in\mathbb{R}^{n_{0}}$, the network computes its output $y_{k}\in\mathbb{R}^{n_{k}}$ as follows:
$y_{0}$ $=u,$
$y_{i+1}$ $=\sigma(W_{i}y_{i}+b_{i}),\quad\text{for all }i\in\{0,1,\ldots,k-1\},$
$N(u)$ $=W_{k}y_{k}+b_{k}.$
Let $y_{i-1}$ and $y_{i}$ denote the input and output of the $i$-th layer, with activation function $\sigma:\to$. The network $N$ is monotone if (i) all weight matrices satisfy $W_{i}\geq 0_{n_{i}\times n_{i+1}}$ and (ii) $\sigma$ is non-decreasing. Monotone neural networks are universal approximators of monotone functions daniels2010monotone .
Using Theorem 3, the search for robust (resp. control) barrier certificates reduces to finding monotone neural networks satisfying conditions (2), (3), and (4) (resp. (5)). However, formally verifying such certificates remains intractable due to the need to evaluate the network over a continuum of states. In the next section, we show how the system’s and certificate’s monotonicity enable safety verification and control synthesis from finite data.
3.2 Monotone Barrier Certificates from Finite Data
In this section, we exploit the system’s and certificates’ monotonicity to develop a data-driven framework for safety verification and control synthesis using only finite data, while ensuring formal correctness. We partition the state space $\mathcal{X}$ and the input space $U$ into hyper-rectangles $\{[\underline{x}^{i},\overline{x}^{i}]\}_{i\in\mathcal{P}}$ and $\{[\underline{u}^{i},\overline{u}^{i}]\}_{i\in\mathcal{S}}$, respectively, where $\mathcal{P}$ and $\mathcal{S}$ are finite index sets corresponding to these partitions. Given $\mathcal{X}_{0}\subseteq\mathcal{X}$ and $\mathcal{X}_{u}\subseteq\mathcal{X}$, we define index sets $\mathcal{I},\mathcal{U}$ as the minimal subsets of $\mathcal{P}$ such that
$\mathcal{X}_{0}\subseteq\bigcup_{i\in\mathcal{I}}[\underline{x}^{i},\overline{x}^{i}],\qquad\mathcal{X}_{u}\subseteq\bigcup_{i\in\mathcal{U}}[\underline{x}^{i},\overline{x}^{i}].$ (6)
By construction, any $x\in\mathcal{X}_{0}$ satisfies $x\leq\overline{x}^{i}$ for some $i\in\mathcal{I}$, and any $x\in\mathcal{X}_{u}$ satisfies $\underline{x}^{i}\leq x$ for some $i\in\mathcal{U}$. For monotone systems with monotone barrier certificates, the points $\{\underline{x}^{i}\}_{i\in\mathcal{U}}$ and $\{\overline{x}^{i}\}_{i\in\mathcal{I}}$ act as worst-case representatives. Thus, the search for robust barrier certificates satisfying (2)-(4) can be conservatively reformulated as finding a monotone neural network $N:\mathcal{X}\to\mathbb{R}$ satisfying:
$N(\overline{x}^{i})$ $\leq\gamma\quad$ $\forall i\in\mathcal{I},$ (7)
$N(\underline{x}^{i})$ $>\eta\quad$ $\forall i\in\mathcal{U},$ (8)
$N(\underline{x}^{i})$ $\leq\delta\quad\implies\quad\max_{j\in\mathcal{S}}N(f(\overline{x}^{i},\overline{u}^{j}))\leq\delta\quad$ $\forall i\in\mathcal{P}.$ (9)
A monotone neural network $N:\mathcal{X}\to$ satisfying (7), (8), and (9) is called a monotone neural robust barrier certificate. Similarly, for safe control synthesis, the search can be reformulated as finding a monotone neural network $N:\mathcal{X}\to\mathbb{R}$ satisfying (7), (8), and
$N(\underline{x}^{i})$ $\leq\delta\quad\implies\quad\min_{j\in\mathcal{S}}N(f(\overline{x}^{i},\underline{u}^{j}))\leq\delta\quad$ $\forall i\in\mathcal{P}.$ (10)
A monotone neural network $N:\mathcal{X}\to$ that satisfies conditions (7), (8), and (9) is referred to as a monotone neural control barrier certificate. Conditions (7)–(10) reduce the verification of neural barrier certificates from the entire (continuous) state set to a finite set of carefully selected maximal and minimal points derived from the coverings of $\mathcal{X}$ and $U$. The following result provides formal guarantees for safety verification and safe control synthesis using monotone neural barrier certificates.
Theorem 4 (Monotone Neural Barrier Certificates).
Consider a monotone dtDS $\mathfrak{S}=(\mathcal{X},\mathcal{X}_{0},U,f)$ with an unknown state transition map $f$ and a set of unsafe states $\mathcal{X}_{u}$. Let $\{[\underline{x}^{i},\overline{x}^{i}]\}_{i\in\mathcal{P}}$ be a family of hyper-rectangles that cover the state set $\mathcal{X}$, $\{[\underline{u}^{i},\overline{u}^{i}]\}_{i\in\mathcal{S}}$ be a family of hyper-rectangles that cover the input set $U$, and the index sets $\mathcal{I},\mathcal{U}$ be the minimal subsets of $\mathcal{P}$ satisfying (6). Then,
  1. 1.
    (Safety verification) If there exists a monotone neural network $N:\mathcal{X}\to$ satisfying (7), (8), and (9), with $\gamma\leq\delta\leq\eta$, then $\mathfrak{S}$ is safe with respect to $\mathcal{X}_{u}$, for every input $\{u_{t}\}_{t=0}^{\infty}\in U$;
  2. 2.
    (Safe control synthesis) If there exists a monotone neural network $N:\mathcal{X}\to$ satisfying (7), (8), and (10), with $\gamma\leq\delta\leq\eta$, then there exists a map $K:\mathcal{X}\to U$ such that $\mathfrak{S}$ with the feedback controller $\{u_{t}=K(x_{t})\}_{t=0}^{\infty}$ is safe with respect to $\mathcal{X}_{u}$.
Proof. We refer to Appendix E for the proof.
Theorem 4 highlights key advantages of our data-driven approach. It allows constructing safety certificates from arbitrary state space partitions, avoiding the curse of dimensionality in fine discretizations anand2023formally and the reliance on SMT solvers for neural network verification edwards2024fossil . By leveraging system monotonicity, it scales to high-dimensional systems and does not require knowledge of Lipschitz constants for the dynamics or barrier function.
3.3 Training Monotone Barrier Certificates.
In this section, we present an adaptive algorithm for training monotone neural robust (resp. control) barrier certificates. Let $\{[\underline{x}^{i},\overline{x}^{i}]\}_{i\in\mathcal{P}}$ and $\{[\underline{u}^{i},\overline{u}^{i}]\}_{i\in\mathcal{S}}$ be hyper-rectangular coverings of the state and input spaces, respectively, and $\mathcal{I},\mathcal{U}$ the index sets defined in (6).
For safety verification, we train a monotone neural network $N:\mathcal{X}\to\mathbb{R}$ as a robust barrier certificate by minimizing the composite mean-squared-error loss $L_{\mathrm{total}}:=L_{1}+L_{2}+L_{3}$, where each term enforces a specific barrier certificate constraint:
$L_{1}$ $:=\mathrm{MSE}(N(\overline{x}^{i}),\gamma),$ $\text{for all }i\in\mathcal{I},$ (11)
$L_{2}$ $:=\mathrm{MSE}(N(\underline{x}^{i}),\eta),$ $\text{for all }i\in\mathcal{U},$ (12)
$L_{3}$ $:=\mathrm{MSE}\left(\min_{j\in\mathcal{S}}N(f(\overline{x}^{i},\underline{u}^{j})),\delta\right),$ $\text{for all }i\in\mathcal{P}.$ (13)
Training uses gradient-based optimization on $L_{\mathrm{total}}$, with tunable parameters $\gamma,\delta,\eta\in\mathbb{R}_{>0}$ satisfying $\gamma<\delta<\eta$. The process stops once $N$ meets conditions (7), (8), and (9).
For safe control synthesis, we jointly train two monotone neural networks: $N:\mathcal{X}\to\mathbb{R}$, which serves as the control barrier certificate, and $K:\mathcal{X}\to\mathcal{U}$, which represents the feedback policy. The objective for both of these networks is the composite mean-squared-error loss $\widehat{L}_{\mathrm{total}}:=L_{1}+L_{2}+L_{4}$, where $L_{1}$ and $L_{2}$ are defined in (11) and (12), and $L_{4}$ is given by
$L_{4}$ $:=\mathrm{MSE}\left(N(f(\overline{x}^{i},K(\overline{x}^{i}))),\delta\right),$ $\forall i\in\mathcal{P}.$ (14)
Training is performed jointly on $N$ and $K$ using gradient-based optimization on $\widehat{L}_{\mathrm{total}}$, with tunable parameters $\gamma,\delta,\eta\in\mathbb{R}_{>0}$ such that $\gamma\leq\delta\leq\eta$. It stops when $N$ satisfies (7), (8), and the following condition:
$N(\underline{x}^{i})\leq\delta\quad\Longrightarrow\quad N(f(\overline{x}^{i},K(\overline{x}^{i})))\leq\delta,\qquad\forall i\in\mathcal{P}.$
$x_{1}$$x_{2}$
(a) Initial Sets
$x_{1}$$x_{2}$
(b) Sampled Points
$x_{1}$$x_{2}$
(c) Refined Sets
Figure 1: Illustration of hyper-rectangular cover refinement for a simple 2D system. The blue and red regions represent the initial and unsafe sets, respectively, while the purple region indicates where condition (10) must be satisfied. Two randomly selected sample points (black dots) are used to update the region where condition (10) is enforced. The condition is then verified at the maximal points of this region, shown as green dots.
Iterative Refinement of the Hyper-rectangular Cover. In practice, the sample complexity for learning monotone neural barrier certificates depends on the granularity of the hyper-rectangular cover $\{[\underline{x}^{i},\overline{x}^{i}]\}_{i\in\mathcal{P}}$ of the state set $\mathcal{X}$. A fine hyper-rectangular cover for $\mathcal{X}$ can make enforcing  (10) across all cells computationally demanding. To address this, we propose an algorithm with linear sample complexity that starts with a coarse hyper-rectangular cover and refines it iteratively as needed to ensure  (10) always holds. Assume that $\mathcal{X}$ is a bounded hyper-rectangle $\mathcal{X}=[\underline{x}_{\min},\overline{x}_{\max}]$ which we consider as a hyper-rectangular cover of $\mathcal{X}$. We also suppose that there exists $\underline{x}_{u}\in\mathcal{X}$ such that $\mathcal{X}_{u}\subseteq[\underline{x}_{u},\overline{x}_{\max}]$. We define the set $T\subseteq\mathcal{X}$ by
$T=[\underline{x}_{u},\overline{x}_{\max}],$ (15)
and compute the set of all maximal points of $H=\mathcal{X}\setminus T$, denoted by $\{\overline{x}^{i}\}_{i\in\mathcal{P}}$ [^2] We refer to Appendix G for details about computing all maximal points of union of hyper-rectangles.. In the first step, our algorithm refines the hyper-rectangular cover $\underline{x}_{\min},\overline{x}_{\max}]$ by replacing it with the family of hyper-rectangles $\{[\underline{x}_{\min},\overline{x}^{i}]\}_{i\in\mathcal{P}}$, and uses this family to define the loss function $L_{\mathrm{total}}$ for training a monotone neural network $N$. If $N$ fails to satisfy condition (10) after a set number of epochs, we sample $m$ intermediate states where the barrier certificate lies between the safe and unsafe thresholds and define the set $Q$ as follows:
$Q:=m\mbox{ samples from }\left\{x\in\mathcal{X}\mid\gamma<N(x)<\eta\right\}.$
Using the sampled points $x\in Q$, we update the set $T$ in (15) as follows:
$T\leftarrow T\cup\bigcup_{x\in Q}[x,\overline{x}_{\max}],$
Then we find all the maximal points of the set $\mathcal{X}\setminus T$ and update the family of hyper-rectangles $\{[\underline{x}_{\min},\overline{x}^{i}]\}_{i\in\mathcal{P}}$  [^3] We refer to Appendix G for details about computing all maximal points of union of hyper-rectangles.. The barrier certificate is then retrained on the refined family of hyper-rectangles, where condition (10) is more likely to hold. The process repeats until all barrier conditions are met over the updated family of hyper-rectangles or a maximum number of iterations is reached. Figure 1 illustrates this iterative update of the family of hyper-rectangles in a 2D example.
Algorithm 1 details the safe control synthesis, and Algorithm 2 covers safety verification. By leveraging monotonicity and carefully designing hyper-rectangular covers of the state space $\mathcal{X}$, our algorithm certifies the safety of unknown monotone systems with linear sample complexity.
Algorithm 1 Monotone Neural Control Barrier Certificate
Input: Simulator access to $f$, sets $\mathcal{X}_{0}$, $\mathcal{X}_{u}$, $\mathcal{X}=[\underline{x}_{\min},\overline{x}_{\max}]$, $e$ total number of epochs, $\hat{e}$ number of epochs before sampling, $m$ number of samples;
Output: Neural networks $N$ (barrier certificate), $K$ (feedback controller);
1:Over-approximate $\mathcal{X}_{0}$, $\mathcal{X}_{u}$, and $\mathcal{X}$ using unions of hyper-rectangles as in (6)
2:Construct maximal initial points $\{\overline{x}^{i}\}_{i\in\mathcal{I}}$ and minimal unsafe points $\{\underline{x}^{i}\}_{i\in\mathcal{U}}$
3:Construct the hyper-rectangular cover $T=[\underline{x}_{u},\overline{x}_{\max}]$
4:Construct $H$ as the set of maximal points of $\mathcal{X}\setminus T$
5:Initialize monotone neural networks NN, $K$ and set hyperparameters $\gamma<\delta=\eta$
6:for $e$ number of epochs do
7:  Minimize loss $\widehat{L}_{\mathrm{total}}=L_{1}+L_{2}+L_{4}$ as defined in (11), (12) and (14), respectively.
8:  if Conditions (7)–(9) not satisfied for $\hat{e}$ epochs then
9:   $Q\leftarrow$ Sample $m$ points from $\{x\in\mathcal{X}\mid\gamma<B(x)<\eta\}$
10:   Update $T\leftarrow T\cup\bigcup_{x\in Q}[x,\overline{x}_{\max}]$
11:   Update $H$ by getting maximal points of $\mathcal{X}\setminus T$
12:   Set $\delta\leftarrow\min_{x\in H}N(x)$
13:  end if
14:  if all constraints are satisfied then
15:   return trained networks NN, $K$
16:  end if
17:end for
18:return Inconclusive (training failed to converge)
Theorem 5 (Correctness and Sample Complexity of Algorithm 1).
Consider a monotone (unknown) dtDS $\mathfrak{S}=(\mathcal{X},\mathcal{X}_{0},U,f)$ with $\mathcal{X}\subseteq\mathbb{R}^{n}$ and an upper closed unsafe set $\mathcal{X}_{u}\subseteq\mathcal{X}$. If Algorithm 1 returns a control barrier certificate $N:\mathcal{X}\to\mathbb{R}$ and a controller $K:\mathcal{X}\to U$, then the algorithm certifies the safety of $\mathfrak{S}$ under the feedback controller $\{u_{t}=K(x_{t})\}_{t=0}^{\infty}$ with respect to the unsafe set $\mathcal{X}_{u}$ using $\mathcal{O}(n)$ simulator queries.
Proof. We refer to Appendix G for the proof.
4 Experiments
This section demonstrates the efficiency of our framework through three case studies, using an RTX 4090 GPU. Both the barrier certificate and feedback controller neural networks consist of $5$ hidden layers with $200$ neurons in each layer and ReLU activation functions. The monotone neural networks are trained using Algorithm 1 with $\hat{e}=1000$ and $e=20000$. While system models are mentioned, they are not used to encode barrier conditions. Table 1 compares training time and sample complexity of Algorithm 1 against state-of-the-art methods, and Table 2 shows scalability of Algorithm 1 for a traffic network from 2 to 100 states. Competing methods either time out after 12 hours or require exponentially many samples, while ours scales efficiently with linear sample growth. Edwards et al. (edwards2024fossil, ) rely on fixed random sampling and SMT-based verification (e.g., Z3 (de2008z3, )), which becomes intractable as dimensions increase.
Table 1: Run time and sample complexity across different methods and case studies. All results are averaged over 30 random initializations. NA indicates failure to converge within 12 hours.
Case Study (size) Edwards et al. edwards2024fossil Anand et al. anand2023formally Zhang et al. zhang2024seev Ours
Running Time
Traffic Control (50D) NA NA NA 20s
Power Grid (13,659D) NA NA NA 250s
Traffic Network (1000D) NA NA NA 135s
Sample Complexity
Traffic Control (50D) 500 $\Omega(2^{50})$ $\Omega(2^{50})$ 51
Power Grid (13,659D) 500 $\Omega(2^{13000})$ $\Omega(2^{13000})$ 13,500
Traffic Network (1000D) 500 $\Omega(2^{1000})$ $\Omega(2^{1000})$ 1001
Table 2: Performance comparison on the Freeway Traffic Network as system dimension grows. Our method scales efficiently, while baselines either fail to converge or require exponential samples. Results are averaged over 30 runs.
Dim Edwards et al. edwards2024fossil Anand et al. anand2023formally Zhang et al. zhang2024seev Ours
Time Samples Time Samples Time Samples Time Samples
2 30s 500 10m $10^{2}$ 12m $10^{2}$ 1s 3
5 50s 500 45m $10^{5}$ 50m $10^{5}$ 2s 6
10 2m 500 NA $\Omega(2^{10})$ NA $\Omega(2^{10})$ 4s 11
25 NA 500 NA $\Omega(2^{25})$ NA $\Omega(2^{25})$ 10s 26
50 NA 500 NA $\Omega(2^{50})$ NA $\Omega(2^{50})$ 20s 51
75 NA 500 NA $\Omega(2^{75})$ NA $\Omega(2^{75})$ 45s 76
100 NA 500 NA $\Omega(2^{100})$ NA $\Omega(2^{100})$ 90s 101
Urban Traffic Control.
Our first case study considers urban traffic control from sadraddini2018formal . The traffic network is modeled as a directed graph $(\mathcal{V},\mathcal{L})$, where each link $l\in\mathcal{L}$ supports one-way flow from tail node $\tau(l)\in\mathcal{V}\cup\{0\}$ to head node $\eta(l)\in\mathcal{V}$, with $\tau(l)=0$ representing external inflows. Our method synthesized a controller in 20s (averaged over 30 runs). Details are in Appendix H.
European Power Grid Verification.
Our second benchmark is the 13659-bus European grid from the pegase project josz2016ac , with lossless swing dynamics discretized as
$\theta_{k+1}=\theta_{k}+\tau\bigl{(}P-BA\sin(B^{\!\top}\theta_{k})\bigr{)}:=f(\theta_{k},P),$ (16)
where $\tau$ is the step size, $\theta\in\mathbb{R}^{n}$ are bus angles, $P\in^{n}$ are power generations/demands at buses, $B$ is the incidence matrix of the interconnection power grid, and $A=\operatorname{diag}(V_{i}V_{j}Y_{ij})$, where $Y_{ij}$ is the line admittances and $V_{i}$ is the voltage amplitude at bus $i$. This power system is a monotone dtDS $\mathfrak{S}=(\mathcal{X},\mathcal{X}_{0},U,f)$ with the state set $\mathcal{X}=[\frac{\pi}{2},\frac{\pi}{2}]^{n}$, the initial set $\mathcal{X}_{0}=[0,0.1]^{n}$, the input set $U=\{P_{\mathrm{nom}}\}$, and the unsafe set $\mathcal{X}_{u}=[\frac{\pi}{4},\frac{\pi}{2}]^{n}$ with $n=13659$, and $\tau=0.1$. Using the Matpower black-box model zimmerman1997matpower , our method certifies safety in 250s on average over 30 runs.
Freeway Traffic Verification.
Finally, we consider a Daganzo cell transmission model with demand and supply constraints, as shown in Figure 4. The traffic network is modeled as a monotone system $\mathfrak{S}=(\mathcal{X},\mathcal{X}_{0},U,f)$ FB-CTDS with the state set $\mathcal{X}=[0,50]^{n}$, the initial set $\mathcal{X}_{0}=[0,5]\times[0,18]\times[0,5]^{n-2}$, the input $U=[0,9]$ and the unsafe set is $\mathcal{X}_{u}=[30,50]^{n}$. Our method finds a monotone neural robust barrier certificate in 135s for a 1000-dimensional network (averaged over 30 runs). See Appendix I for details.
5 Conclusions
This work advances neurosymbolic verification by proposing a scalable framework for synthesizing safety controllers in high-dimensional monotone systems using monotone neural barrier certificates. Leveraging the systems’ order-preserving structure, our approach reduces verification to local boundary checks, achieving potentially linear sample complexity without explicit models or conservative assumptions. Experiments on large-scale benchmarks demonstrate its practicality and scalability. Future work may extend the framework to systems with partial or mixed monotonicity and explore integration with reinforcement learning for adaptive control.
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: The claims match theoretical and experimental results.
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: We have explicitly mentioned that this work only supports monotone systems.
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory assumptions and proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [Yes]
  14. Justification: Every theorem is backed by its proof (if novel) or referenced.
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental result reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: We have included architecture of neural networks, dynamical systems, and hyper-parameters. Code is also available.
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [Yes]
  24. Justification: Code is available.
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental setting/details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: Every detail is mentioned in experiments’ section.
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment statistical significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [No]
  34. Justification: We have formal proof, which is deterministic.
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments compute resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [Yes]
  39. Justification: We trained on a single RTX 4090, as mentioned in the paper.
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code of ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Justification: We have faithfully followed Code of Ethics.
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [Yes]
  49. Justification: We have disclosed our broader impact. Our work paves the way for AI agents to be implemented in real-life scenarios with safety guarantees.
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [N/A]
  54. Justification: Our method ensures safety.
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [Yes]
  59. Justification: We have followed the guidelines.
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [Yes]
  64. Justification: We have included every detail about our models and training.
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and research with human subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [N/A]
  69. Justification: Paper does not involve crowdsourcing nor research with human subjects.
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional review board (IRB) approvals or equivalent for research with human subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [N/A]
  74. Justification: The paper does not involve crowdsourcing nor research with human subjects.
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
  76. 16.
    Declaration of LLM usage
  77. Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research? Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
  78. Answer: [N/A]
  79. Justification: LLMs were used only for editing.
  80. Guidelines:
    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.
Appendix A Related Work
Barrier Certificates. [38] introduced barrier certificates (BCs) as functions whose level sets over-approximate system reachable sets. Control barrier certificates (CBCs) later emerged as a key tool for synthesizing safe controllers [3, 49, 24, 11, 15], typically using Sum-of-Squares (SOS) optimization [39, 57, 44].
Neural Barrier Certificates. [55] first introduced neural network–based barrier certificates, using a shallow network and MILP for formal verification. Later, [35] employed SMT solvers to find counterexamples for iterative refinement. Neural barrier certificates have since been extended to hybrid systems [56] and stochastic dynamics [32].
Neural Control Barrier Certificates. Neural control barrier certificates (NCBCs) extend these ideas to safe controller synthesis. Recent work [41, 29, 18, 17, 31, 30] learns both barrier functions and control policies using neural networks. Verification typically relies on SMT solvers [1, 55, 20], reachable set methods [48], polynomial approximations [45], Lipschitz bounds [5], or ReLU verification [26]. Recently, [54, 53] an exact NCBC verification algorithm was proposed that synthesizes a controller if the barrier is valid. Most of these methods, however, assume full system knowledge (except [5]) and rely heavily on SMT solvers, which scale poorly to deep networks due to their exponential complexity.
Monotone dynamics.
Monotone systems are dynamical systems whose trajectories preserve a partial order, exhibiting regular behaviors common in applications like traffic networks [14] and biological systems [47]. While monotonicity has been shown to reduce sample complexity in Lyapunov analysis [40], its role in constructing barrier certificates remains less explored.
Appendix B Proof of Theorem 1
Appendix C Proof of Theorem 2
Appendix D Proof of Theorem 3
Before proving Theorem 3 part 1, we introduce several notations and intermediate results that will be used throughout the proof. Consider a monotone dtDS $\mathfrak{S}=(\mathcal{X},\mathcal{X}_{0},U,f)$. The reachable set of $\mathfrak{S}$, denoted by $\Omega\subseteq\mathcal{X}$, is defined as follows:
$\Omega$ $=\bigcup_{T\in\mathbb{Z}_{\geq 0}}\Omega_{T},$ (17)
[…]
We first state the following intermediate result about the reachable set $\Omega$.
Lemma 1.
Consider a monotone dtDS $\mathfrak{S}=(\mathcal{X},\mathcal{X}_{0},U,f)$ with the reachable set $\Omega$ defined in (17). Then, for every $x\in\Omega$ and every $u\in U$, we have $f(x,u)\in\Omega$.
Proof.
Let $x\in\Omega$ and $u\in U$. Then there exists $T\in\mathbb{Z}_{\geq 0}$ such that $x\in\Omega_{T}$. This means that there exists $\{x_{t}\}_{t=0}^{\infty}$ a trajectory of $\mathfrak{S}$ with the input $\{u_{t}\}_{t=0}^{\infty}$ starting from $x_{0}\in\mathcal{X}_{0}$ such that $x=x_{T}$. Now consider the input $\{w_{t}\}_{t=0}^{\infty}$ defined by
$w_{t}=\begin{cases}u_{t}&t\leq T,\\ u&t>T.\end{cases}$ (18)
Let $\{y_{t}\}_{t=0}^{\infty}$ be the trajectory of $\mathfrak{S}$ with the input $\{w_{t}\}_{t=0}^{\infty}$ starting from $x_{0}$. By definition of the input $\{w_{t}\}_{t=0}^{\infty}$ in (18), it is clear that $y_{t}=x_{t}$, for every $t\leq T$. Moreover, for $t=T+1$, we have
$y_{T+1}=f(y_{T},w_{T})=f(y_{T},u)=f(x_{T},u)=f(x,u),$
where the second equality holds by definition of the input $\{w_{t}\}_{t=0}^{\infty}$ in (18), the third equality holds using the fact that $x_{t}=y_{t}$, for every $t\leq T$, and the fourth equality holds because $x=x_{T}$. Therefore, by definition of $\Omega_{T+1}$, we have $f(x,u)\in\Omega_{T+1}$ and thus $f(x,u)\in\Omega$. ∎
Motivated by the definition of the reachable set of $\mathfrak{S}$ in (17) we define the set $\widehat{\Omega}\subseteq\mathcal{X}$ as the smallest lower closed set in $\mathcal{X}$ containing $\Omega$. More specifically, $\widehat{\Omega}$ is defined as:
$\widehat{\Omega}=\{x\in\mathcal{X}\;|\;x\leq y,\mbox{ for some }y\in\Omega\}.$ (19)
Given $\gamma,\eta\in_{\geq 0}$ such that $\gamma\leq\eta$, we use the lower closed set $\widehat{\Omega}$ to define the function $B:\mathcal{X}\to$ as follows:
$B(x)=\begin{cases}\gamma&x\in\widehat{\Omega}\\ \eta+1&x\not\in\widehat{\Omega}\end{cases}$ (20)
The next lemma proves a useful property of the function $B$.
Lemma 2.
Consider a monotone dtDS $\mathfrak{S}=(\mathcal{X},\mathcal{X}_{0},U,f)$ with the lower closed set $\widehat{\Omega}$ as defined in (19). Then the function $B:\mathcal{X}\to$ defined in (20) is monotone.
Proof.
Let $x,y\in\mathcal{X}$ be such that $x\leq y$. We consider two cases i) $y\in\widehat{\Omega}$ and ii) $y\not\in\widehat{\Omega}$.
  1. i)
    In this case, by definition of $\widehat{\Omega}$ in (19), there exists a point $z\in\Omega$ such that $y\leq z$. Moreover, we have $x\leq y\leq z$. Therefore, we have $x\in\widehat{\Omega}$ and $B(x)=\gamma=B(y)$.
  2. ii)
    In this case, we have $B(y)=\eta+1$. Since we know that $\gamma\leq\eta<\eta+1$, we get $B(x)\leq\eta+1=B(y)$.
Thus, in both cases i) and ii), we have $B(x)\leq B(y)$. This implies that $B$ is a monotone map. ∎
Now we go back to the proof of Theorem 3 part 1. We divide the proof into if part and only if part.
Proof of if part.
Suppose that $B:\mathcal{X}\to$ is a monotone function satisfying conditions (2),(3), and (4). By definition 2, $B:\mathcal{X}\to$ is a robust barrier certificate for $\mathfrak{S}$ with respect to $\mathcal{X}_{u}$. Therefore, by Theorem 1, $\mathfrak{S}$ is safe with respect to $\mathcal{X}_{u}$, for every input $\{u_{t}\}_{t=0}^{\infty}$.
Proof of only if part.
Now assume that the monotone dtDS $\mathfrak{S}$ is safe with respect to upper closed unsafe set $\mathcal{X}_{u}$, for every input $\{u_{t}\}_{t=0}^{\infty}\in U$. Let $\widehat{\Omega}$ be the lower closed set defined in (19) and $B:\mathcal{X}\to$ be the function defined in (20). Our goal is to show that $B:\mathcal{X}\to$ defined in (20) is a monotone robust barrier certificate for $\mathfrak{S}$ satisfying conditions (2),(3), and (4). First, by Lemma 2, the function $B:\mathcal{X}\to$ is monotone.
  1. 1.
    ($B$ satisfies condition (2)) Suppose that $x\in\mathcal{X}_{0}$. By the definition of $\Omega_{0}$, this implies that $x\in\Omega_{0}$. As a result, we have $x\in\Omega\subseteq\widehat{\Omega}$. Using the definition of $B$ in (20), we get $B(x)=\gamma$. This shows that $B$ satisfies condition (2).
  2. 2.
    ($B$ satisfies condition (3)) Suppose that $x\in\mathcal{X}_{u}$. Since $\mathfrak{S}$ is safe with respect to $\mathcal{X}_{u}$ for any input $\{u_{t}\}_{t=0}^{\infty}\in U$, we get that $x\not\in\Omega$. Note that $\widehat{\Omega}$ is the smallest lower closed set containing $\Omega$ and $\mathcal{X}_{u}$ is an upper closed set. As a result, $x\not\in\widehat{\Omega}$. By definition of $B$ in (20), we have $B(x)=\eta+1>\eta$. This shows that $B$ satisfies condition (3).
  3. 3.
    ($B$ satisfies condition (4)) Let $x\in\mathcal{X}$ be such that $B(x)\leq\delta$. Note that $\delta\leq\eta<\eta+1$. By definition of the function $B$ in (20), this implies that $x\in\widehat{\Omega}$. By definition of the set $\widehat{\Omega}$ in (19), there exists $y\in\Omega$ such that $x\leq y$. Since the state transition map $f$ is monotone, for every $u\in U$, we have $f(x,u)\leq f(y,u)$. Note that $y\in\Omega$ and $u\in U$. Therefore, using Lemma 1, we can show that $f(y,u)\in\Omega$. Using the definition of the lower closed set $\widehat{\Omega}$ in (19), this implies that $f(x,u)\in\widehat{\Omega}$. As a result, one can use the definition of $B$ in (20) to show that $B(f(x,u))=\gamma$. Since this equality holds, for every $u\in U$, we get
    $\sup_{u\in U}B(f(x,u))=\gamma\leq\delta.$
    This shows that $B$ satisfies condition (4).
This completes the proof of Theorem 3 part 1.
Now we focus on the proof of Theorem 3 part 2. Before we state the proof, we introduce several notations and intermediate results that will be used throughout the proof. Consider a monotone dtDS $\mathfrak{S}=(\mathcal{X},\mathcal{X}_{0},U,f)$ with a map $K:\mathcal{X}\to U$. The reachable set of $\mathfrak{S}$ with the feedback controller $u=K(x)$, which is denoted by $\Omega^{K}\subseteq\mathcal{X}$, is defined by
$\Omega^{K}$ $=\bigcup_{T\in\mathbb{Z}_{\geq 0}}\Omega^{K}_{T},$ (21)
[…]
Motivated by the definition of the reachable set of $\mathfrak{S}$ in (17) we define the set $\widehat{\Omega}^{K}\subseteq\mathcal{X}$ as the smallest lower closed set in $\mathcal{X}$ containing $\Omega^{K}$. More specifically, $\widehat{\Omega}^{K}$ is defined as:
$\widehat{\Omega}^{K}=\{x\in\mathcal{X}\;|\;x\leq y,\mbox{ for some }y\in\Omega^{K}\}.$ (22)
Given $\gamma,\eta\in_{\geq 0}$ such that $\gamma\leq\eta$, we use the lower closed set $\widehat{\Omega}^{K}$ to define the function $B^{K}:\mathcal{X}\to$ as follows:
$B^{K}(x)=\begin{cases}\gamma&x\in\widehat{\Omega}^{K}\\ \eta+1&x\not\in\widehat{\Omega}^{K}\end{cases}$ (23)
The next lemma proves a useful property of the function $B^{K}$.
Lemma 3.
Consider a monotone dtDS $\mathfrak{S}=(\mathcal{X},\mathcal{X}_{0},U,f)$ with the lower closed set $\widehat{\Omega}$ as defined in (19). Then the function $B^{K}:\mathcal{X}\to$ defined in (23) is monotone.
Proof.
Let $x,y\in\mathcal{X}$ be such that $x\leq y$. We consider two cases: i) $y\in\widehat{\Omega}^{K}$, and ii) $y\not\in\widehat{\Omega}^{K}$.
  1. 1.
    In this case, there exists $z\in\Omega^{K}$ such that $y\leq z$. Note that we have $x\leq y\leq z$. This implies that $x\in\widehat{\Omega}^{K}$. As a result, by definition of $B^{K}$ in (23). we have $B^{K}(x)=B^{K}(y)=\gamma$.
  2. 2.
    In this case, by definition of $B^{K}$ in (23), we have $B^{K}(y)=\eta$. This implies that $B^{K}(y)=\eta+1$, since Note that $\gamma\leq\eta<\eta+1$. This means that $B(x)\leq\eta+1=B(y)$.
In both cases i) and ii), we have $B^{K}(x)\leq B^{K}(y)$. This implies that $B^{K}$ is a monotone map. ∎
Now we get back to Theorem 3 part 2. We divide the proof into if part and only if part.
Proof of if part.
Suppose that $B:\mathcal{X}\to$ is a monotone barrier certificate satisfying (2),(3), and (5). Then, by Theorem 2, there exists a map $K:\mathcal{X}\to$ such that the system $\mathfrak{S}$ with the feedback controller $\{u_{t}=K(x_{t})\}_{t=0}^{\infty}$ is safe with respect to $\mathcal{X}_{u}$.
Proof of only if part.
Suppose that there exists a map $K:\mathcal{X}\to$ such that the system $\mathfrak{S}$ with the feedback controller $\{u_{t}=K(x_{t})\}_{t=0}^{\infty}$ is safe with respect to $\mathcal{X}_{u}$. Let the lower closed set $\widehat{\Omega}^{K}$ be defined in (19) and the function $B^{K}:\mathcal{X}\to$ be defined as in 23. Our goal is to show that $B^{K}:\mathcal{X}\to$ is a monotone control barrier certificate for $\mathfrak{S}$ satisfying conditions (2),(3), and (5). By Lemma (3), it is clear that $B^{K}$ is monotone.
  1. 1.
    ($B^{K}$ satisfies condition (2)) Suppose that $x\in\mathcal{X}_{0}$. By the definition of $\Omega^{K}_{0}$, this implies that $x\in\Omega^{K}_{0}$. As a result, we have $x\in\Omega^{K}\subseteq\widehat{\Omega}^{K}$. Using the definition of $B^{K}$ in (23), we get $B^{K}(x)=\gamma$. This shows that $B^{K}$ satisfies condition (2).
  2. 2.
    ($B^{K}$ satisfies condition (3)) Suppose that $x\in\mathcal{X}_{u}$. Since $\mathfrak{S}$ with the feedback controller $u=K(x)$ is safe with respect to $\mathcal{X}_{u}$, we get that $x\not\in\Omega^{K}$. Note that $\widehat{\Omega}^{K}$ is the smallest lower closed set containing $\Omega^{K}$ and $\mathcal{X}_{u}$ is an upper closed set. As a result, $x\not\in\widehat{\Omega}^{K}$. By definition of $B^{K}$ in (23), we have $B^{K}(x)=\eta+1>\eta$. This shows that $B^{K}$ satisfies condition (3).
  3. 3.
    ($B^{K}$ satisfies condition (5)) Let $x\in\mathcal{X}$ be such that $B^{K}(x)\leq\delta$. Note that $\delta\leq\eta<\eta+1$. By definition of the function $B^{K}$ in (23), this implies that $x\in\widehat{\Omega}^{K}$. By definition of the set $\widehat{\Omega}^{K}$ in (22), there exists $y\in\Omega^{K}$ such that $x\leq y$. Since the state transition map $f$ is monotone, for every $u\in U$, we have $f(x,K(y))\leq f(y,K(y))$. Note that $y\in\Omega^{K}$. Therefore, by definition of $\Omega^{K}$ in (21), we have $f(y,K(y))\in\Omega^{K}$. Using the definition of the lower closed set $\widehat{\Omega}^{K}$ in (19), this implies that $f(x,K(y))\in\widehat{\Omega}$. As a result, we have
    $\inf_{u\in U}B^{K}(f(x,u))\leq B^{K}(f(x,K(y)))=\gamma\leq\delta,$
    where the first inequality holds by monotonicity of $f$ and $B^{K}$ and the second equality holds by definition of $B^{K}$ in (20). This shows that $B^{K}$ satisfies condition (5).
This completes the proof of Theorem 3 part 2.
Appendix E Proof of Theorem 4
Regarding part 1, our goal is to show that the monotone neural network $N:\mathcal{X}\to$ satisfying conditions (7), (8), and (9) is a robust barrier certificate for $\mathfrak{S}$ with respect to the unsafe set $\mathcal{X}_{u}$. More specifically, our goal is to show that the monotone neural network $N:\mathcal{X}\to$ satisfies conditions (2), (3), and (4).
  1. 1.
    (NN satisfies condition (2)) Let $x\in\mathcal{X}_{0}$. Since $\mathcal{X}_{0}\subseteq\bigcup_{i\in\mathcal{I}}[\underline{x}^{i},\overline{x}^{i}]$, there exists $k\in\mathcal{I}$ such that $x\leq\overline{x}^{k}$. Since $N:\mathcal{X}\to$ satisfies condition (7), we have $N(\overline{x}^{k})\leq\gamma$. On the other hand, NN is a monotone neural network, thus $N(x)\leq N(\overline{x}^{k})\leq\gamma$. This shows that the monotone neural network $N$ satisfies condition (2).
  2. 2.
    (NN satisfies condition (3)) Let $x\in\mathcal{X}_{u}$. Since $\mathcal{X}_{u}\subseteq\bigcup_{i\in\mathcal{U}}[\underline{x}^{i},\overline{x}^{i}]$, there exists $j\in\mathcal{U}$ such that $\underline{x}^{j}\leq x$. Since $N$ satisfies the condition (8), we have $\eta\leq N(\underline{x}^{j})$. Using the monotonicity of the neural network NN, we get $\eta\leq N(\underline{x}^{j})\leq N(x)$. This means that the monotone neural network $N$ satisfies condition (3).
  3. 3.
    (NN satisfies condition (4)) Let $x\in\mathcal{X}$ be such that $N(x)\leq\delta$. Since the family of hyper-rectangles $\{[\underline{x}^{i},\overline{x}^{i}]\}_{i\in\mathcal{P}}$ covers the state set $\mathcal{X}$, there exists $k\in\mathcal{P}$ such that $x\in[\underline{x}^{k},\overline{x}^{k}]$. Using the monotonicity of $N$, we get $N(\underline{x}^{k})\leq N(x)\leq\delta$. Since $N$ satisfies condition (9), we have $\max_{j\in\mathcal{S}}N(f(\overline{x}^{k},\overline{u}^{j}))\leq\delta$. As a result, for every $u\in U$, we have
    $N(f(x,u))\leq N(f(\overline{x}^{k},u))\leq\max_{j\in\mathcal{S}}N(f(\overline{x}^{k},\overline{u}^{j}))\leq\delta,$
    where the first inequality holds by the monotonicity of $N$ and $f$ and the fact that $x\leq\overline{x}^{k}$, the second inequality holds by monotonicity of $N$ and $f$ and the fact that $u\leq\max_{j\in\mathcal{S}}\overline{u}^{j}$. This means that $N$ satisfies condition (4).
Therefore, the monotone neural network $N:\mathcal{X}\to$ satisfies conditions (2), (3), and (4). The result then follows using Theorem (1). This completes the proof of part 1.
Regarding part 2, our goal is to show that the monotone neural network $N:\mathcal{X}\to$ satisfying conditions (7), (8), and (10) is a control barrier certificate for $\mathfrak{S}$ with respect to $\mathcal{X}_{u}$. More specifically, our goal is to show that $N$ satisfies conditions (2), (3), and (5).
  1. 1.
    (NN satisfies condition (2)) Let $x\in\mathcal{X}_{0}$. Since $\mathcal{X}_{0}\subseteq\bigcup_{i\in\mathcal{I}}[\underline{x}^{i},\overline{x}^{i}]$, there exists $k\in\mathcal{I}$ such that $x\leq\overline{x}^{k}$. Since $N:\mathcal{X}\to$ satisfies condition (7), we have $N(\overline{x}^{k})\leq\gamma$. On the other hand, NN is a monotone neural network, thus $N(x)\leq N(\overline{x}^{k})\leq\gamma$. This shows that the monotone neural network $N$ satisfies condition (2).
  2. 2.
    (NN satisfies condition (3)) Let $x\in\mathcal{X}_{u}$. Since $\mathcal{X}_{u}\subseteq\bigcup_{i\in\mathcal{U}}[\underline{x}^{i},\overline{x}^{i}]$, there exists $j\in\mathcal{U}$ such that $\underline{x}^{j}\leq x$. Since $N$ satisfies the condition (8), we have $\eta\leq N(\underline{x}^{j})$. Using the monotonicity of the neural network NN, we get $\eta\leq N(\underline{x}^{j})\leq N(x)$. This means that the monotone neural network $N$ satisfies condition (3).
  3. 3.
    (NN satisfies condition (5)) Let $x\in\mathcal{X}$ be such that $N(x)\leq\delta$. Since the family of hyper-rectangles $\{[\underline{x}^{i},\overline{x}^{i}]\}_{i\in\mathcal{P}}$ covers the state set $\mathcal{X}$, there exists $k\in\mathcal{P}$ such that $x\in[\underline{x}^{k},\overline{x}^{k}]$. Using the monotonicity of $N$, we get $N(\underline{x}^{k})\leq N(x)\leq\delta$. Since $N$ satisfies condition (10), we have $\min_{j\in\mathcal{S}}N(f(\overline{x}^{k},\overline{u}^{j}))\leq\delta$. Since both $N$ and $f$ are monotone functions, for every $u\in U$, we have $N(f(x,u))\leq N(f(\overline{x}^{k},u))$. By taking the infimum of both sides of this inequality, we obtain
    $\inf_{u\in U}N(f(x,u))\leq\inf_{u\in U}N(f(\overline{x}^{k},u))\leq\min_{j\in\mathcal{S}}N(f(\overline{x}^{k},\overline{u}^{j}))\leq\delta,$ (24)
    where the second inequality holds using the fact that $\{[\underline{u}^{i},\overline{u}^{i}]\}_{i\in\mathcal{S}}$ is a hyper-rectangular cover for $U$. The inequality (24) implies that $N$ satisfies condition (5).
Therefore, the monotone neural network $N:\mathcal{X}\to$ satisfies conditions (2), (3), and (5). The result then follows using Theorem (2). This completes the proof of part 2.
Appendix F Algorithm for Safety Verification
In this appendix, we present an algorithm, analogous to Algorithm (1), for constructing monotone neural robust barrier certificates.
Algorithm 2 Monotone Neural Robust Barrier Certificate
Input: Simulator access to $f$, sets $\mathcal{X}_{0}$, $\mathcal{X}_{u}$, $\mathcal{X}=[\underline{x}_{\min},\overline{x}_{\max}]$, $e$ total number of epochs, $\hat{e}$ number of epochs before sampling, $m$ number of samples;
Output: Neural network $B$ (barrier certificate);
1:Over-approximate $\mathcal{X}_{0}$, $\mathcal{X}_{u}$, and $\mathcal{X}$ using unions of hyper-rectangles as in (6)
2:Construct maximal initial points $\{\overline{x}^{i}\}_{i\in\mathcal{I}}$ and minimal unsafe points $\{\underline{x}^{i}\}_{i\in\mathcal{U}}$
3:Construct the hyper-rectangular cover $T=[\underline{x}_{u},\overline{x}_{\max}]$
4:Construct $H$ as the set of maximal points of $\mathcal{X}\setminus T$
5:Initialize monotone neural networks $B$ and set hyperparameters $\gamma<\delta=\eta$
6:for $e$ number of epochs do
7:  Minimize loss $\widehat{L}_{\mathrm{total}}=L_{1}+L_{2}+L_{4}$ as defined in (11)- (13), respectively.
8:  if Conditions (7)–(9) not satisfied for $\hat{e}$ epochs then
9:   $Q\leftarrow$ Sample $m$ points from $\{x\in\mathcal{X}\mid\gamma<B(x)<\eta\}$
10:   Update $T\leftarrow T\cup\bigcup_{x\in Q}[x,\overline{x}_{\max}]$
11:   Update $H$ by getting maximal points of $\mathcal{X}\setminus T$
12:   Set $\delta\leftarrow\min_{x\in H}N(x)$
13:  end if
14:  if all constraints are satisfied then
15:   return trained network $B$
16:  end if
17:end for
18:return Inconclusive (training failed to converge)
Appendix G Correctness and Sample Complexity of Algorithm 1
In this appendix, we provide additional details and proofs for Algorithm (1) presented in Section (3.3). We begin by presenting a systematic approach for computing the maximal elements of the set $\mathcal{X}\setminus T$.
G.1 Computing Maximal Points of $\mathcal{X}\setminus T$
[…]
$T=\bigcup_{x\in Q\cup\{\underline{x}_{u}\}}[x,\overline{x}_{\max}].$ (25)
Then, we can characterize the set $\mathcal{X}\setminus T$ as follows:
$\mathcal{X}\setminus T=\left\{y\in\mathcal{X}\,\middle|\,\forall x\in Q\cup\{\underline{x}_{u}\},\,\exists i\in\{1,\ldots,n\}\text{ such that }y_{i}<x_{i}\right\}.$ (26)
That is, a point $y\in\mathcal{X}$ lies in the region $\mathcal{X}\setminus T$ if it is strictly smaller than every $x\in Q\cup\{\underline{x}_{u}\}$ in at least one coordinate.
To conservatively cover $\mathcal{X}\setminus T$, we construct a set $\mathcal{H}$ of maximal points defining hyper-rectangular regions. This is done through a structured procedure as follows:
Step 1: Incomparable base sampling. We begin by sampling a set $Q=\{y^{1},\ldots,y^{m}\}\subset\mathcal{X}\setminus T$ such that all points are pairwise incomparable. These serve as the base points for constructing $\mathcal{H}$.
Step 2: Maximal combinations. For each nonempty subset $S\subseteq\{y^{1},\ldots,y^{m}\}$, we compute the element-wise maximum
$x_{S}:=\max\{y\mid y\in S\},$
and collect the resulting points into a set $\mathcal{H}_{1}$. These represent the upper envelope of the base points.
Step 3: Anchor points. To ensure full coverage of the domain boundaries, we define a set of anchor points. For each coordinate $i\in\{1,\ldots,n\}$, define $z^{i}\in\mathcal{X}$ by
$(z^{i})_{j}=\begin{cases}(\overline{x}_{\max})_{j}&j\neq i\\ \min_{y\in Q}y_{i}&j=i\end{cases}$
We collect the resulting points in the set $\mathcal{H}_{2}$, for every $i\in\{1,\ldots,n\}$.
Step 4: Dominance filtering. We define the final set $\mathcal{H}$ as
$\mathcal{H}:=\text{clipmax}(\mathcal{H}_{1}\cup\mathcal{H}_{2}),$ (27)
where $\text{clipmax}(\cdot)$ denotes the subset of points not dominated by any other in the union. That is, we remove all $x\in\mathcal{H}_{1}\cup\mathcal{H}_{2}$ such that there exists $x^{\prime}\neq x$ with $x^{\prime}\leq x$.
We will need the following Lemma to show that $\mathcal{H}$ consists of maximal points of the set $\mathcal{X}\setminus T$.
Lemma 4 (Complement of a hyper-rectangle).
Let $a\in\mathcal{X}$ and $A=[a,\overline{x}_{\max}]$ be a hyper-rectangle. Then we have
$\mathcal{X}\setminus A\subseteq\bigcup_{y\in\mathcal{G}}[\underline{x}_{\min},y],$
where $\mathcal{G}=\{a\}\cup\bigcup_{i=1}^{n}\{v^{i}\}$, where for every $i\in\{1,\ldots,n\}$, the point $v^{i}$ is defined as follows:
$v^{i}_{j}=\begin{cases}(\overline{x}_{\max})_{j}&j\neq i\\ a_{i}&j=i\end{cases}$
Proof.
Let $w\in\mathcal{X}\setminus A$. This means that $w\not\geq a$. As a result, there exists $i\in\{1,\ldots,n\}$ such that $w_{i}<a_{i}$. Now consider the point $v^{i}$, for every $j\neq i$, we have $v^{i}_{j}=(\overline{x}_{\max})_{j}\geq w_{j}$. Moreover, we have $v^{i}_{i}=a_{i}>w_{i}$. This means that $v^{i}>w$ and therefore $w\in[\underline{x}_{\min},v^{i}]$. ∎
Now we extend the previous result for two hyper-rectangles.
Lemma 5 (Complement of union of two hyper-rectangles).
Let $a,b\in\mathcal{X}$ and $A=[a,\overline{x}_{\max}],B=[b,\overline{x}_{\max}]$ be two hyper-rectangles. Then we have
$\mathcal{X}\setminus(A\cup B)\subseteq\bigcup_{y\in\mathcal{G}}[\underline{x}_{\min},y],$
where $\mathcal{G}=\{x_{S}\}\cup\bigcup_{i=1}^{n}\{v^{i}\}$, where $x_{S}$ is element-wise maximum of $a$ and $b$ and for every $i\in\{1,\ldots,n\}$, the point $v^{i}$ is defined as follows:
$v^{i}_{j}=\begin{cases}(\overline{x}_{\max})_{j}&j\neq i\\ \min\{a_{i},b_{i}\}&j=i\end{cases}$
Proof.
We know that $\mathcal{X}\setminus(A\cup B)=(\mathcal{X}\setminus A)\bigcap(\mathcal{X}\setminus B)$. By Lemma 4, we have
$\mathcal{X}\setminus A\subseteq\bigcup_{y\in\mathcal{G}}[\underline{x}_{\min},y],\qquad\qquad\mathcal{X}\setminus B\subseteq\bigcup_{\widehat{y}\in\mathcal{\widehat{G}}}[\underline{x}_{\min},\widehat{y}$],
where $\mathcal{G}=\{a\}\cup\bigcup_{i=1}^{n}\{v^{i}\}$, where for every $i\in\{1,\ldots,n\}$, the point $v^{i}$ is defined as follows:
$v^{i}_{j}=\begin{cases}(\overline{x}_{\max})_{j}&j\neq i\\ a_{i}&j=i\end{cases}$
and $\mathcal{\widehat{G}}=\{b\}\cup\bigcup_{i=1}^{n}\{\widehat{v}^{i}\}$, where for every $i\in\{1,\ldots,n\}$, the point $\widehat{v}^{i}$ is defined as follows:
$\widehat{v}^{i}_{j}=\begin{cases}(\overline{x}_{\max})_{j}&j\neq i\\ b_{i}&j=i\end{cases}$
Therefore, we have
[…]
$=\left([\underline{x}_{\min},a]\cup\bigcup_{i\in\{1,\ldots,n\}}[\underline{x}_{\min},v^{i}\}]\right)\bigcap\left([\underline{x}_{\min},b]\cup\bigcup_{i\in\{1,\ldots,n\}}[\underline{x}_{\min},\widehat{v}^{i}\}]\right)$
$\subseteq\left([\underline{x}_{\min},x_{S}]\cup\bigcup_{i\in\{1,\ldots,n\}}[\underline{x}_{\min},v^{i}\}]\right)\bigcap\left([\underline{x}_{\min},x_{S}]\cup\bigcup_{i\in\{1,\ldots,n\}}[\underline{x}_{\min},\widehat{v}^{i}\}]\right)$
$\subseteq[\underline{x}_{\min},x_{S}]\cup\left(\bigcup_{i\in\{1,\ldots,n\}}[\underline{x}_{\min},\min\{v^{i},\widehat{v}^{i}\}]\right),$
where the first equality follows from Lemma 4, the third inequality follows from the interval inclusions $\underline{x}_{\min},a]\subseteq[\underline{x}_{\min},x_{S}]$ and $\underline{x}_{\min},b]\subseteq[\underline{x}_{\min},x_{S}]$, and the last inequality follows from the distributive property of unions and intersections. ∎
Lemma 5 can be extended to an arbitrary number of hyper-rectangles by applying mathematical induction.
Theorem 6 (Maximal Points of $\mathcal{X}\setminus T$).
Let $\mathcal{X}\subseteq\mathbb{R}^{n}$, and let $T=\bigcup_{x\in Q\cup\{\underline{x}_{u}\}}[x,\overline{x}_{\max}]$ be a finite union of upper hyper-rectangles. Suppose the set $\mathcal{H}$ is constructed through the procedure in Steps 1–4, as described in (27). Then
$\mathcal{X}\setminus T\subseteq\bigcup_{y\in\mathcal{H}}[\underline{x}_{\min},y].$
Proof.
The results follows by applying Lemma 5 inductively. ∎
Theorem 7 (Cardinality of $\mathcal{H}$).
Let $\mathcal{X}\subseteq\mathbb{R}^{n}$, and let $T=\bigcup_{x\in Q\cup\{\underline{x}_{u}\}}[x,\overline{x}_{\max}]$ be a finite union of upper hyper-rectangles. Suppose the set $\mathcal{H}$ is constructed through the procedure in Steps 1–4, as described in (27). Then, the cardinality of the set $\mathcal{H}$ is $\mathcal{O}(n)$.
Proof.
By the procedure in Steps 1–4, we have $\mathcal{H}\subseteq\mathcal{H}_{1}\cup\mathcal{H}_{2}$. This implies that $|\mathcal{H}|\leq|\mathcal{H}_{1}|+|\mathcal{H}_{2}|$. Based on Step 2, we have $|\mathcal{H}_{1}|\leq 2^{m}-1$. Moreover, in Step 3, we construct $n$ anchor points. This means that $|\mathcal{H}_{2}|=n$. As a result, the cardinality of $\mathcal{H}$ is at most $2^{m}+n-1$ and therefore $\mathcal{O}(n)$. ∎
G.2 Proof of Theorem 5
Proof.
We divide the proof into two parts: (i) proof of correctness, and (ii) proof of linear sample complexity.
  1. 1.
    (Correctness). Suppose that Algorithm 1 converges and returns a monotone neural control barrier certificate $N$ and a feedback controller $K$. By the termination condition of Algorithm 1, $N$ must satisfy conditions (7)–(8). We now show that $N$ and $K$ also satisfy condition (9). Since the set $H$ consists of the maximal elements of $\mathcal{X}\setminus T$, and $\delta=\min_{x\in H}N(x)$, it follows that for every $x\in T$, we have $N(x)\geq\delta$. Therefore, condition (9) holds trivially on the region $T$ for the monotone neural control barrier certificate $N$ and the feedback controller $K$. Furthermore, the algorithm ensures that condition (9) also holds on $\mathcal{X}\setminus T$. Thus, $N$ and $K$ satisfy condition (9) over the entire state space $\mathcal{X}$. The result then follows directly from Theorem 4.
  2. 2.
    (Linear sample complexity). The sample complexity of Algorithm (1) is exactly the number of elements in the family of hyper-rectangles $\{[\underline{x}_{\min},\overline{x}^{i}]_{i\in\mathcal{P}}\}$ where $\overline{x}^{i}$ are the maximal points of the set $\mathcal{X}\setminus T$. By Theorem 7, the cardinality of the maximal points of $\mathcal{X}\setminus T$ is $|\mathcal{H}|$ and it is of $\mathcal{O}(n)$. This completes the proof of linear sample complexity.
Appendix H Urban Traffic Control
We consider an urban traffic network described by a family of links (roads) categorized according to their actuation type: $\mathcal{L}_{r}$ (roads with traffic light control), $\mathcal{L}_{o}$ (on-ramps controlled by ramp meters), and $\mathcal{L}_{f}$ (freeway segments without direct control). Uncontrolled roads are also treated like freeways, so the set of all links (roads) in this network is given by $\mathcal{L}=\mathcal{L}_{r}\cup\mathcal{L}_{o}\cup\mathcal{L}_{f}$. The number of vehicles on link $l$ at time $t$ is $x_{[l],t}\in[0,c_{l}]$, where $c_{l}$ is the capacity of link $l$. We assume that the traffic flow follows fluid-like behavior. Here, the state space is $\mathcal{X}:=\Pi_{l\in\mathcal{L}}[0,c_{l}]$. The outflow from link $l$ at time $t$ is:
$q_{[l],t}=\min\Big{\{}x_{[l],t},\;\bar{q}_{l},\;\min_{l^{\prime}:\tau(l^{\prime})=\eta(l)}\frac{\alpha_{l:l^{\prime}}}{\beta_{l:l^{\prime}}}(c_{l^{\prime}}-x_{[l^{\prime}],t})\Big{\}},$ (28)
where $\bar{q}_{l}$ is max outflow from $l$, $\alpha_{l:l^{\prime}}$ is capacity ratio of link $l^{\prime}$ relative to $l$, and $\beta_{l:l^{\prime}}$ is turning ratio from $l$ to $l^{\prime}$. Figure 2 presents a schematic diagram of the link outflow $q[l]$ as a function of the state variables $x[l]$ and $x[l^{\prime}]$, commonly referred to as the fundamental diagram in traffic flow literature [23]. The flow $q[l]$ decreases when one or more downstream links lack sufficient capacity to accommodate incoming vehicles. In such cases—when the last term in (28) determines the minimum—we say the traffic is congested. Otherwise, if the upstream demand governs the flow, the traffic is considered to be in the free-flow regime. We define the congestion-free set $\Pi$ as:
$\Pi:=\left\{x\in\mathcal{X}\mid\min\{x_{[l]},\bar{q}_{l}\}\leq\frac{\alpha_{l:l^{\prime}}}{\beta_{l:l^{\prime}}}(c_{l^{\prime}}-x_{[l^{\prime}]})\;\forall l,l^{\prime}:\tau(l^{\prime})=\eta(l)\right\}.$ (29)
The actual allowable outflow from link $l$ depends on its actuation:
$\tilde{q}_{[l],t}=\begin{cases}s_{[l],t}\cdot q_{[l],t}&\text{if }l\in\mathcal{L}_{r},\\ \min\{q_{[l],t},r_{[l],t}\}&\text{if }l\in\mathcal{L}_{o},\\ q_{[l],t}&\text{if }l\in\mathcal{L}_{f}.\end{cases}$ (30)
Here, $s_{[l],t}\in\{0,1\}$ indicates the traffic light status (1 for green, 0 for red), and $r_{[l],t}\in\mathbb{R}_{+}$ is the ramp metering rate, both of which are control input to the system. Thus, $U=\mathbb{R}_{>0}^{4}\times\{0,1\}^{14}$. To prevent conflicting greens, for any pair $l,l^{\prime}$ with perpendicular intersections, we require $s_{[l],t}+s_{[l^{\prime}],t}\leq 1$. The state transition map $f$ for this urban traffic network is defined by:
$x_{[l],t+1}=x_{[l],t}-\tilde{q}_{[l],t}+\sum_{l^{\prime}:\tau(l)=\eta(l^{\prime})}\beta_{l^{\prime}:l}\tilde{q}_{[l^{\prime}],t}:=f_{[l]}(x_{t},s_{t},r_{t}),\qquad\mbox{ for all }l\in\mathcal{L}.$ (31)
We define the set of unsafe states as $\mathcal{X}_{u}=\mathcal{X}\setminus\Pi$; that is, a state is considered unsafe if at least one link in the network is congested. Some trajectories of the system are depicted in Figure 3. We refer the reader to [42] for more details on model of this system, its topology, and proof of monotonicty. Parameters of the model are found in Table 3
$x_{[l]},x_{[l^{\prime}]}$$q_{[l]}$$\frac{\alpha_{l:l^{\prime}}}{\beta_{l:l^{\prime}}}(c_{l^{\prime}}-x_{[l^{\prime}]})$free flowcongested flow
Figure 2: Fundamental diagram. The outflow from link $l$ decreases when the number of vehicles on its immediate downstream link $l^{\prime}$ approaches the downstream capacity. This reduction in flow, caused by limited downstream supply, characterizes the onset of congestion due to blocking behavior.
Figure 3: Trajectories of the system. Dotted lines indicate entering congestion status.
Table 3: Parameters of the urban traffic network
links parameters
1–26, 49–53 $\bar{q}=15$, $c=40$
27–34 $\bar{q}=15$, $c=30$
35–48 $\bar{q}=40$, $c=60$
Turning ratios value
$\beta_{2:50},\beta_{4,53},\beta_{8:51},\beta_{12:7},\beta_{13:28},\beta_{15:30},\beta_{16:28}$ 0.2
$\beta_{21:32},\beta_{24:32},\beta_{26:2},\beta_{36:31},\beta_{36:33},\beta_{39:27},\beta_{43:29}$
$\beta_{5:12},\beta_{6:13},\beta_{6:18},\beta_{10:21},\beta_{10:26}$ 0.3
$\beta_{1:20},\beta_{6:7}$ 0.4
$\beta_{1:2},\beta_{11:12},\beta_{14:30},\beta_{17:7},\beta_{17:18},\beta_{19:2},\beta_{19:20}$
$\beta_{31:22},\beta_{31:25},\beta_{33:24},\beta_{49:3},\beta_{49:50},\beta_{51:4},\beta_{52:5}$ 0.5
$\beta_{22:34},\beta_{23:24},\beta_{23:34},\beta_{27:14},\beta_{27:17},\beta_{29:16},\beta_{52:53}$
$\beta_{2:3},\beta_{3:4},\beta_{4:5},\beta_{8:9},\beta_{12:13},\beta_{13:14},\beta_{15:16},\beta_{16:17}$
$\beta_{20:21},\beta_{21:22},\beta_{24:25},\beta_{25:26},\beta_{36:37},\beta_{39:40}$ 0.8
$\beta_{43:44},\beta_{46:47}$
Capacity ratios value
$\alpha_{19:2},\alpha_{26:2},\alpha_{17:7},\alpha_{12:7},\alpha_{13:28},\alpha_{16:28}$ $0.5$
$\alpha_{14:30},\alpha_{15:30},\alpha_{21:32},\alpha_{24:32},\alpha_{22:34},\alpha_{23:34}$
Appendix I Freeway Traffic Verification
For the freeway traffic network $\mathfrak{S}=(\mathcal{X},\mathcal{X}_{0},U,f)$ shown in Figure (4), the state space is given by $\mathcal{X}=[0,50]^{n}$, and the state transition map $f:\mathcal{X}\times U\to\mathcal{X}$ is defined by:
$f_{1}(q)$ $=q_{1}+\tau\left(F_{0\rightarrow 1}-F_{1\rightarrow 2}(q)\right),$
$f_{i}(q)$ $=q_{i}+\tau\left(F_{i-1\rightarrow i}(q)-F_{i\rightarrow i+1}(q)\right),\quad i\in\{2,\ldots,n-1\},$
$f_{n}(q)$ $=q_{n}+\tau\left(F_{n-1\rightarrow n}(q)-F_{n\rightarrow 0}(q)\right),$ (32)
where the input is the on-ramp inflow $F_{0\rightarrow 1}\in U=[0,9]$, the parameter $\tau=0.1$ is sampling time, and $q=\begin{bmatrix}q_{1}&\ldots&q_{n}\end{bmatrix}^{\top}$ with $q_{i}\geq 0$ representing the vehicle density in road segment $i$. The flow from road segment $i$ to $j$, denoted $F_{i\rightarrow j}(q)$, is defined as:
$F_{i\rightarrow j}(q)$ $=\min\left\{\varphi_{i}(q_{i}),\sigma_{j}(q_{j})\right\},$ (33)
$F_{n\rightarrow 0}(q)$ $=\varphi_{n}(q_{n}),$ (34)
where $\varphi_{i}(q_{i})$ models the increasing outflow demand of segment $i$, and $\sigma_{j}(q_{j})$ models the decreasing inflow supply of segment $j$. The demand and supply functions are given by:
$\varphi_{i}(q_{i})$ $=15\left(1-e^{-0.038q_{i}}\right),\quad\sigma_{i}(q_{i})=\mathrm{ReLU}\left(15\left(1-\left(\frac{q_{i}}{100}\right)^{4}\right)\right),\quad\forall i\neq 2,$
$\varphi_{2}(q_{2})$ $=10\left(1-e^{-0.138q_{2}}\right),\quad\sigma_{2}(q_{2})=\mathrm{ReLU}\left(10\left(1-\left(\frac{q_{2}}{50}\right)^{4}\right)\right).$ (35)
The initial set for this system is defined as $\mathcal{X}_{0}=[0,5]\times[0,18]\times[0,5]^{n-2}$. The unsafe set is $\mathcal{X}_{u}=[30,50]^{n}$, which indicates congestion in all roads. We refer the reader to [9] for more details on this system and proof of monotonicity. For $n=1000$, our Algorithm (F) find a monotone neural robust barrier certificate for $\mathfrak{S}$ in 135s (averaged over 30 runs).
Figure 4: A line digraph representing an $n$-dim traffic network.

### v3

@@
 Monotone Neural Barrier Certificates
-Abstract
-This work presents a neurosymbolic framework for safety verification and control synthesis in high-dimensional monotone dynamical systems, using an algorithm with linear sample complexity—without relying on explicit models or conservative Lipschitz bounds.
+Abstract.
+This report presents a neurosymbolic framework for safety verification and control synthesis in high-dimensional monotone dynamical systems without relying on explicit models or conservative Lipschitz bounds.
 The approach combines the expressiveness of neural networks with the rigor of symbolic reasoning via barrier certificates, functional analogs of inductive invariants that formally guarantee safety.
@@
 This enables scalable, formally sound verification directly from simulation data, bridging black-box learning and formal guarantees within a unified neurosymbolic framework.
-Empirical results on three large-scale benchmarks—a freeway traffic model with increasing size, a 50-dimensional urban traffic network, and a 13,000-dimensional power grid—demonstrate the scalability and effectiveness of the approach in real-world, safety-critical systems.
-1 Introduction
+1. Introduction
 Discrete-time dynamical systems offer a foundational framework for modeling and controlling a broad spectrum of digitally controlled autonomous systems, including autonomous vehicles, implantable medical devices, biological networks, and industrial process management.
 As these systems constitute a critical component of our safety infrastructure, high-assurance formal reasoning is essential to ensure their safe operation.
-A principled and efficient approach to providing such safety guarantees involves synthesizing barrier certificates prajna2006barrier ; ADA-SC-ME-GN-KS-PT:19 —a functional analog of inductive invariants—whose existence serves as a formal proof of safety AP-AR:04 .
-Although finding such certificates is inherently computationally challenging, recent neuro-symbolic approaches have shown considerable promise through the use of neural barrier certificates chang2019neural ; zhao2020synthesizing ; zhang2024exact ; robey2020learning .
+A principled and efficient approach to providing such safety guarantees involves synthesizing barrier certificates [24, 1]—a functional analog of inductive invariants—whose existence serves as a formal proof of safety [23].
+Although finding such certificates is inherently computationally challenging, recent neuro-symbolic approaches have shown considerable promise through the use of neural barrier certificates [7, 35, 34, 27].
 While these neural certificates provide substantial flexibility in searching for safety guarantees, we posit that scalability and sample efficiency can be further improved in specialized control applications by identifying and leveraging structural properties of the underlying system.
-In this work, we focus on a subclass of discrete-time dynamical systems known as monotone systems angeli2003monotone ; HLS:95 , which frequently arise in domains such as queuing networks farina2011positive , population dynamics leenheer2004predator , biological systems DA:13 , and chemical networks EDS:07 .
+In this work, we focus on a subclass of discrete-time dynamical systems known as monotone systems [5, 29], which frequently arise in domains such as queuing networks [15], population dynamics [18], biological systems [4], and chemical networks [30].
 By leveraging their structural characteristics, we synthesize barrier certificates represented as monotone neural networks, which we refer to as monotone neural barrier certificates.
 Monotone Systems.
-Monotone dynamical systems HLS:95 ; angeli2003monotone are characterized by the property that their trajectories preserve a natural partial order induced on their state set.
-As a paradigmatic example, consider traffic flow dynamics sadraddini2018formal , where increasing the number of vehicles on a road segment does not lead to a decrease in vehicle density downstream under normal flow conditions.
-That is, if one traffic configuration contains more vehicles than another, the evolution of the system preserves this ordering.
-Similar examples arise in domains such as biochemical reactions in synthetic biology EDS:07 , frequency regulation in power system, congestion control in transportation networks SC-MA:15 ; GC-EL-KS:15 , population dynamics in ecology, and compartmental models in epidemiology angeli2003monotone .
-Related concepts have also been studied in other domains, including positive linear systems farina2011positive , well-structured transition systems FINKEL200163 , and well-behaved transition systems Blondin2016WellBT .
-It is well known that monotone systems exhibit highly ordered transient and asymptotic behaviors HLS:95 , making them particularly well-suited for data-driven methods, as their inherent structure helps mitigate the challenges posed by limited or noisy data.
-While monotonicity has informed control and stability analysis AR:15 ; SC:19 , its potential for learning-based safety verification remains largely untapped.
+Monotone dynamical systems [29, 5] are characterized by the property that their trajectories preserve a natural partial order induced on their state set.
+Examples of monotone systems arise in domains such as biochemical reactions in synthetic biology [30], frequency regulation in power grids, congestion control in transportation networks [10, 8], population dynamics in ecology, and compartmental models in epidemiology [5].
+Related concepts have also been studied in other domains, including positive linear systems [15], well-structured transition systems [16], and well-behaved transition systems [6].
+It is well known that monotone systems exhibit highly ordered transient and asymptotic behaviors [29], making them particularly well-suited for data-driven methods, as their inherent structure helps mitigate the challenges posed by limited or noisy data.
+While monotonicity has informed control and stability analysis [26, 9], its potential for learning-based safety verification remains largely untapped.
 Barrier Certificates (A Proof of Safety).
@@
 A common alternative is to construct an inductive state-invariant, i.e., a set that contains the initial states and is closed under the system’s transitions.
-Barrier certificates prajna2006barrier offer one such approach by defining a real-valued function over the state space whose level sets form a “barrier” separating safe and unsafe regions.
+Barrier certificates [24] offer one such approach by defining a real-valued function over the state space whose level sets form a “barrier” separating safe and unsafe regions.
 Specifically, the barrier function must: 1) take values less than or equal to a constant over the initial states, 2) and strictly greater than that constant over unsafe states; and 3) decrease along system trajectories outside the unsafe set.
@@
 Despite their intuitive appeal and ease of verification, complexity of finding CBCs remains a significant hurdle in their widespread adoption.
-Traditional methods for synthesizing CBCs rely on optimization-based formulations such as Sum-of-squares (SOS) programming AP-SJ:05 ; kong2013exponential .
+Traditional methods for synthesizing CBCs rely on optimization-based formulations such as Sum-of-squares (SOS) programming [21, 17].
 SOS is a popular approach that is applicable to systems with polynomial dynamics; however, it quickly becomes intractable due to the growth of semidefinite constraints with system dimension and polynomial degree.
-Linear programming (LP) methods, while more tractable, are typically limited to linear or affine certificates and cannot capture the rich nonlinear boundaries present in many real-world systems SS-XC-EA:13 ; YZ-etal:16 .
-Recent progress in deep learning has enabled the use of neural networks to be trained as barrier certificates anand2023formally ; lindemann2024learning ; dawson2022safe ; liu2023safe ; zhao2020synthesizing ; Rus2023 ; zhang2024exact .
+Linear programming (LP) methods, while more tractable, are typically limited to linear or affine certificates and cannot capture the rich nonlinear boundaries present in many real-world systems [28, 32].
+Recent progress in deep learning has enabled the use of neural networks to be trained as barrier certificates [3, 19, 12, 20, 35, 31, 34].
 These so-called neural control barrier certificates (NCBCs) represent both the barrier function and controller as neural networks.
-While expressive, these approaches often require full knowledge of the system dynamics (to verify correctness using constraint satisfaction tools such as Z3 de2008z3 ) or rely on bounding Lipschitz constants (which entails exponential sample complexity), limiting scalability and applicability to black-box systems anand2023formally ; zhang2024exact .
-Moreover, these approaches treat the system as unstructured, ignoring properties like monotonicity that could drastically reduce the number of required samples and constraints [^1] We refer to Appendix A for a comprehensive review of related works..
+While expressive, these approaches often require full knowledge of the system dynamics (to verify correctness using constraint satisfaction tools such as Z3 [13]) or rely on bounding Lipschitz constants (which entails exponential sample complexity), limiting scalability and applicability to black-box systems [3, 34].
+Moreover, these approaches treat the system as unstructured, ignoring properties like monotonicity that could drastically reduce the number of required samples and constraints.
 Monotone Neural Control-Barrier Certificates.
-We introduce a machine learning algorithm with linear sample complexity for synthesizing control barrier certificates specifically tailored to (unknown) monotone systems.
+We introduce an efficient scheme for synthesizing barrier certificates specifically tailored to (unknown) monotone systems.
 By exploiting monotonicity, we restrict the satisfaction of safety conditions to critical boundary slices of the state space—namely, the upper boundary of the initial set and the lower boundary of the unsafe set.
 This reduces the verification burden from the entire state space to a carefully selected subset—without compromising formal guarantees—and enables the use of gradient-based optimization to learn barrier certificates using simulation data alone.
-Our method avoids the need for explicit models, Lipschitz bounds, or semidefinite relaxations, and yields an algorithm for synthesizing barrier certificates that scales linearly with the system dimension.
+Our method avoids the need for explicit models, Lipschitz bounds, or semidefinite relaxations.
 This makes it uniquely suited for safety verification in large-scale, nonlinear systems where monotonicity arises naturally from physical or structural properties.
-Contributions.
-In summary, this work introduces a learning framework for safety verification and safe control synthesis of unknown monotone dynamical systems using monotone neural barrier certificates.
-The key contributions of our paper are as follows:
-  1. 1.
-    Monotone structure of barrier certificates.
-We investigate the structure of barrier certificates for monotone dynamical systems and establish a theoretical result demonstrating that, under mild conditions on the unsafe set, monotone barrier certificates can effectively characterize the system’s safety.
-This result shows that restricting to monotone certificates incurs no loss of generality and supports their use to improve sample complexity and computational cost of safety verification.
-  2. 2.
-    Monotonicity as a catalyst for scalable verification.
-By leveraging the monotone structure of barrier certificates, we propose a scalable framework that verifies safety using only a finite amount of simulator data—without requiring explicit models or Lipschitz constants—by reducing verification to localized boundary checks.
-  3. 3.
-    Monotone Neural Barrier Certificates.
-We propose an algorithm that employs monotone neural networks to learn barrier certificates by leveraging reduced safety conditions on the boundaries of the state set.
-Unlike existing methods based on sum-of-squares, linear programming, or neural barrier certificates, our algorithm achieves linear sample complexity—rather than exponential—and supports arbitrary partitioning of the state set.
-  4. 4.
-    Validation on large-scale, safety-critical domains.
-We evaluate the efficiency of our algorithm on three challenging case studies: (i) a 1,000-dimensional freeway traffic model, (ii) a 50-dimensional urban traffic control network, and (iii) a 13,000-dimensional power grid.
-These results demonstrate the practicality and scalability of our method for safety-critical real-world applications.
-Moreover, in the high-dimensional freeway traffic model, we show that our algorithm outperforms state-of-the-art methods for synthesizing barrier certificates.
-2 Preliminaries and Mathematical Foundations
+2. Preliminaries and Mathematical Foundations
 Notations.
@@
 A hyper-rectangle is denoted by $\underline{x},\overline{x}]:=\{x\in{\mathbb{R}}^{n}\mid\underline{x}\leq x\leq\overline{x}\}$.
-Given a set $A\subseteq^{n}$, $x\in^{n}$ is a maximal (resp.
+Given a set $A\subseteq\mathbb{R}^{n}$, $x\in\mathbb{R}^{n}$ is a maximal (resp.
 minimal) point of $A$, if there does not exists any $y\in A$ such that $x\leq y$ (resp.
 $y\leq x$).
-2.1 Safety of Dynamical Systems
+2.1. Safety of Dynamical Systems
 In this section, we provide a formal definition of discrete-time dynamical systems and establish a rigorous notion of safety. We then present two central safety objectives.
-Definition 1 (Discrete-time dynamical system).
+Definition 2.1 (Discrete-time dynamical system).
 A discrete-time control system (dtDS) is a tuple $\mathfrak{S}:=(\mathcal{X},\mathcal{X}_{0},U,f)$, where $\mathcal{X}\subseteq{\mathbb{R}}^{n}$ represents the state set, $\mathcal{X}_{0}\subseteq\mathcal{X}$ is the initial state set, $U\subseteq{\mathbb{R}}^{m}$ is the set of inputs, and $f:\mathcal{X}\times U\to\mathcal{X}$ is the state transition map.
 Given an input sequence $\{u_{t}\}_{t=0}^{\infty}\in U$, the evolution of $\mathfrak{S}$ under this input is described by:
-$x_{t+1}=f(x_{t},u_{t}),\qquad\text{for all }t\geq 0.$ (1)
-The sequence $\{x_{t}\}_{t=0}^{\infty}$ defined by (1) is called a trajectory of $\mathfrak{S}$ for the input $\{u_{t}\}_{t=0}^{\infty}$.
+(2.1) $x_{t+1}=f(x_{t},u_{t}),\qquad\text{for all }t\geq 0.$
+The sequence $\{x_{t}\}_{t=0}^{\infty}$ defined by (2.1) is called a trajectory of $\mathfrak{S}$ for the input $\{u_{t}\}_{t=0}^{\infty}$.
 Given a dtDS $\mathfrak{S}=(\mathcal{X},\mathcal{X}_{0},U,f)$, a set of unsafe states $\mathcal{X}_{u}\subseteq\mathcal{X}$, and an input sequence $\{u_{t}\}_{t=0}^{\infty}$, we say that $\mathfrak{S}$ is safe with respect to $\mathcal{X}_{u}$ for this input sequence if all trajectories starting from initial states in $\mathcal{X}_{0}$ remain outside of $\mathcal{X}_{u}$.
@@
 In safety verification, inputs are viewed as disturbances, and the goal is to guarantee system safety under all possible input sequences $\{u_{t}\}_{t=0}^{\infty}\in U$.
-To this end, robust barrier certificates have been widely employed as an effective tool for safety verification (prajna, ).
-Definition 2 (Robust barrier certificates).
+To this end, robust barrier certificates have been widely employed as an effective tool for safety verification [25].
+Definition 2.2 (Robust barrier certificates).
 Consider a dtDS $\mathfrak{S}=(\mathcal{X},\mathcal{X}_{0},U,f)$ with a set of unsafe states $\mathcal{X}_{u}\subseteq\mathcal{X}$.
 A function $B:\mathcal{X}\rightarrow{\mathbb{R}}$ is called a robust barrier certificate (RBC) for $\mathfrak{S}$ with respect to the unsafe set $\mathcal{X}_{u}\subseteq\mathcal{X}$ if there exists $\gamma,\delta,\eta\in{\mathbb{R}}_{\geq 0}$ such that $\gamma\leq\delta\leq\eta$, and:
-$B(x)\leq\gamma,\qquad\text{ for all }x\in\mathcal{X}_{0},$ (2)
-$B(x)>\eta,\qquad\text{ for all }x\in\mathcal{X}_{u},\text{ and }$ (3)
-$B(x)\leq\delta\implies\mathrm{sup}_{u\in U}\{B(f(x,u))\}\leq\delta.$ (4)
-The following theorem formulates the safety verification problem as a search for RBCs prajna .
-Theorem 1 (Safety verification via robust barrier certificates).
+(2.2) $B(x)\leq\gamma,\qquad\text{ for all }x\in\mathcal{X}_{0},$
+(2.3) $B(x)>\eta,\qquad\text{ for all }x\in\mathcal{X}_{u},\text{ and }$
+(2.4) $B(x)\leq\delta\implies\mathrm{sup}_{u\in U}\{B(f(x,u))\}\leq\delta.$
+The following theorem formulates the safety verification problem as a search for RBCs [25].
+Theorem 2.3 (Safety verification via robust barrier certificates).
 Consider a dtDS $\mathfrak{S}=(\mathcal{X},\mathcal{X}_{0},U,f)$ and a set of unsafe states $\mathcal{X}_{u}\subseteq\mathcal{X}$.
-Suppose that there exists a robust barrier certificate satisfying conditions (2)-(4).
+Suppose that there exists a robust barrier certificate satisfying conditions (2.2)-(2.4).
 For any input sequence $\{u_{t}\}_{t=0}^{\infty}\in U$, the system $\mathfrak{S}$ is safe with respect to $\mathcal{X}_{u}$.
-Proof. We refer to Appendix B for the proof.
 Safe Control Synthesis Problem.
 In safe control synthesis, inputs are treated as control policies, aiming to synthesize a controller $\{u_{t}\}_{t=0}^{\infty}$ that ensures system safety.
-Control barrier certificates have been widely adopted as an effective tool for synthesizing such controllers (ADA-XX-JWG-PT:17, ).
-Definition 3 (Control barrier certificates).
+Control barrier certificates have been widely adopted as an effective tool for synthesizing such controllers [2].
+Definition 2.4 (Control barrier certificates).
 Consider a dtDS $\mathfrak{S}=(\mathcal{X},\mathcal{X}_{0},U,f)$ with a set of unsafe states $\mathcal{X}_{u}\subseteq\mathcal{X}$.
-A function $B:\mathcal{X}\rightarrow\mathbb{R}$ is called a Control Barrier Certificate (CBC) for $\mathfrak{S}$ with respect to the unsafe set $\mathcal{X}_{u}\subseteq\mathcal{X}$ if there exists $\gamma,\delta,\eta\in\mathbb{R}_{\geq 0}$ such that $\gamma\leq\delta\leq\eta$ and $B$ satisfy conditions (2), (3), and the following condition
-$B(x)\leq\delta\implies\mathrm{inf}_{u\in U}\{B(f(x,u))\}\leq\delta.$ (5)
-The next result formulates the safe control synthesis problem as a search for CBCs ADA-XX-JWG-PT:17 .
-Theorem 2 (Control synthesis via control barrier certificates).
+A function $B:\mathcal{X}\rightarrow{\mathbb{R}}$ is called a Control Barrier Certificate (CBC) for $\mathfrak{S}$ with respect to the unsafe set $\mathcal{X}_{u}\subseteq\mathcal{X}$ if there exists $\gamma,\delta,\eta\in{\mathbb{R}}_{\geq 0}$ such that $\gamma\leq\delta\leq\eta$ and $B$ satisfy conditions (2.2), (2.3), and the following condition
+(2.5) $B(x)\leq\delta\implies\mathrm{inf}_{u\in U}\{B(f(x,u))\}\leq\delta.$
+The next result formulates the safe control synthesis problem as a search for CBCs [2].
+Theorem 2.5 (Control synthesis via control barrier certificates).
 Consider a dtDS $\mathfrak{S}=(\mathcal{X},\mathcal{X}_{0},U,f)$ and a set of unsafe states $\mathcal{X}_{u}\subseteq\mathcal{X}$.
-Suppose that there exists a control barrier certificate satisfying conditions (2), (3), and (5).
+Suppose that there exists a control barrier certificate satisfying conditions (2.2), (2.3), and (2.5).
 Then, there exists a map $K:\mathcal{X}\to U$ such that the system $\mathfrak{S}$ with the feedback controller $\{u_{t}=K(x_{t})\}_{t=0}^{\infty}$ is safe with respect to $\mathcal{X}_{u}$.
-Proof. We refer to Appendix C for the proof.
-2.2 Monotone Dynamical Systems
+2.2. Monotone Dynamical Systems
 In this subsection, we introduce a special class of discrete-time dynamical systems known as monotone systems. We begin by formally defining the notion of monotonicity.
-Definition 4 (Monotone maps).
+Definition 2.6 (Monotone maps).
 Consider the sets $\mathcal{X}{\subseteq}{\mathbb{R}}^{n}$ and $\mathcal{Y}{\subseteq}{\mathbb{R}}^{m}$.
@@
 Using this notion of monotonicity, we define monotone dynamical systems.
-Definition 5 (Monotone dynamical systems).
+Definition 2.7 (Monotone dynamical systems).
 A dtDS $\mathfrak{S}=(\mathcal{X},\mathcal{X}_{0},U,f)$ is monotone if the state transition map $f:\mathcal{X}\times U\to\mathcal{X}$ is monotone.
-For a differentiable transition map $f$ and convex sets $\mathcal{X}$ and $U$, the system $\mathfrak{S}$ is monotone if and only if $Df(x,u)\geq 0_{n\times(n+m)}$ for all $(x,u)\in\mathcal{X}\times U$ HLS:95 .
-Monotone systems preserve the standard partial order along their trajectories HLS:95 ; angeli2003monotone , resulting in highly regular behaviors that facilitate reasoning about dynamics via partial orders.
+For a differentiable transition map $f$ and convex sets $\mathcal{X}$ and $U$, the system $\mathfrak{S}$ is monotone if and only if $Df(x,u)\geq 0_{n\times(n+m)}$ for all $(x,u)\in\mathcal{X}\times U$ [29].
+Monotone systems preserve the standard partial order along their trajectories [29, 5], resulting in highly regular behaviors that facilitate reasoning about dynamics via partial orders.
 Problem Formulation.
@@
 While $f$ is not explicitly known, we assume access to a simulator that returns $f(x,u)$ for any $x\in\mathcal{X}$ and $u\in U$.
-This paper focuses on the following problem.
-Problem 1 (Safety Certificates for Unknown Monotone Systems).
+This report focuses on the following problem.
+Problem 2.8 (Safety Certificates for Unknown Monotone Systems).
 Consider a monotone dtDS $\mathfrak{S}=(\mathcal{X},\mathcal{X}_{0},U,f)$ with an unknown state-transition map $f$, and a set of unsafe states $\mathcal{X}_{u}\subseteq\mathcal{X}$.
 Our goal is to use a finite set of samples from a simulator of $f$ to:
-  1. 1.
+  1. (1)
     (Safety verification): Verify that $\mathfrak{S}$ is safe with respect to $\mathcal{X}_{u}$ for all input $\{u_{t}\}_{t=0}^{\infty}\in U$.
-  2. 2.
+  2. (2)
     (Safe control synthesis): Design a feedback controller $K:\mathcal{X}\rightarrow U$ such that $\mathfrak{S}$ with the feedback controller $\{u_{t}=K(x_{t})\}_{t=0}^{\infty}$ is safe with respect to $\mathcal{X}_{u}$.
-The monotonicity assumption in Problem 1, although made in the absence of an explicit model, is often justified by prior knowledge of the system’s physical properties.
+The monotonicity assumption in Problem 2.8, although made in the absence of an explicit model, is often justified by prior knowledge of the system’s physical properties.
 Many physical processes naturally exhibit monotonicity, where increases in inputs or states lead to non-decreasing responses.
-This property is common in domains such as traffic networks SC-MA:15 , power systems park2020uniqueness , biological processes EDS:07 , and supply chains yu2019equilibrium .
-3 Monotone Neural Barrier Certificates
+This property is common in domains such as traffic networks [10], power systems [22], biological processes [30], and supply chains [33].
+3. Monotone Neural Barrier Certificates
 We propose a data-driven framework that leverages system monotonicity to learn barrier functions using neural networks.
 This approach enables the construction of sound safety certificates from limited data, avoiding full system identification or exhaustive sampling.
-We first analyze the structure of barrier certificates for monotone systems and show that, under certain assumptions on the unsafe set $\mathcal{X}_{u}$, any safe monotone system admits a monotone robust (resp.
-control) barrier certificate.
-3.1 Monotone Barrier Certificates
-In this subsection, we analyze the structure of barrier certificates for safety verification and control synthesis in a monotone system $\mathfrak{S}=(\mathcal{X},\mathcal{X}_{0},U,f)$.
-We show that if the unsafe set $\mathcal{X}_{u}$ is upper closed, both problems can be addressed using monotone barrier certificates.
-Theorem 3 (Monotone Barrier Certificates).
-Consider a monotone dtDs $\mathfrak{S}:=(\mathcal{X},\mathcal{X}_{0},U,f)$ with an upper-closed set of unsafe states $\mathcal{X}_{u}\subseteq\mathcal{X}$. Then, the following statements hold:
-  1. 1.
+3.1. Monotone Barrier Certificates
+In this subsection, we propose a suitable structure of barrier certificates that enables efficient data-driven safety verification and control synthesis for the monotone system $\mathfrak{S}=(\mathcal{X},\mathcal{X}_{0},U,f)$.
+Theorem 3.1 (Monotone Barrier Certificates).
+Consider a monotone dtDs $\mathfrak{S}:=(\mathcal{X},\mathcal{X}_{0},U,f)$ with an unsafe states $\mathcal{X}_{u}\subseteq\mathcal{X}$. Then, the following statements hold:
+  1. (1)
     (Safety verification).
-$\mathfrak{S}$ is safe with respect to $\mathcal{X}_{u}$, for every input $\{u_{t}\}_{t=0}^{\infty}\in U$ if and only if there exists a monotone robust barrier certificate satisfying (2),(3), and (4).
-  2. 2.
+$\mathfrak{S}$ is safe with respect to $\mathcal{X}_{u}$, for every input $\{u_{t}\}_{t=0}^{\infty}\in U$ if there exists a monotone robust barrier certificate satisfying (2.2),(2.3), and (2.4).
+  2. (2)
     (Safe control synthesis).
-there exists a map $K:\mathcal{X}\to U$ such that $\mathfrak{S}$ with the feedback controller $\{u_{t}=K(x_{t})\}_{t=0}^{\infty}$ is safe with respect to $\mathcal{X}_{u}$ if and only if there exists a monotone control barrier certificate satisfying (2),(3), and (5).
-Proof. We refer to Appendix D for the proof.
-For monotone systems, Theorem 3 reduces the search for robust (control) barrier certificates to monotone barrier functions.
+there exists a map $K:\mathcal{X}\to U$ such that $\mathfrak{S}$ with the feedback controller $\{u_{t}=K(x_{t})\}_{t=0}^{\infty}$ is safe with respect to $\mathcal{X}_{u}$ if there exists a monotone control barrier certificate satisfying (2.2),(2.3), and (2.5).
+Proof. We refer to Appendix A for the proof.
+For monotone systems, Theorem 3.1 provides a sufficient condition for searching for robust (control) barrier certificates within the class of monotone barrier functions.
 As will be shown in Section 3.2, the monotonicity structure of barrier certificates plays a crucial role in our framework, leading to significantly lower sample complexity and computational cost—particularly in high-dimensional systems.
@@
 $N(u)$ $=W_{k}y_{k}+b_{k}.$
-Let $y_{i-1}$ and $y_{i}$ denote the input and output of the $i$-th layer, with activation function $\sigma:\to$.
+Let $y_{i-1}$ and $y_{i}$ denote the input and output of the $i$-th layer, with activation function $\sigma:\mathbb{R}\to\mathbb{R}$.
 The network $N$ is monotone if (i) all weight matrices satisfy $W_{i}\geq 0_{n_{i}\times n_{i+1}}$ and (ii) $\sigma$ is non-decreasing.
-Monotone neural networks are universal approximators of monotone functions daniels2010monotone .
-Using Theorem 3, the search for robust (resp.
-control) barrier certificates reduces to finding monotone neural networks satisfying conditions (2), (3), and (4) (resp.
-(5)).
+Monotone neural networks are universal approximators of monotone functions [11].
+By Theorem 3.1, the search for robust (respectively, control) barrier certificates reduces to finding monotone neural networks that satisfy conditions (2.2), (2.3), and (2.4) (respectively, (2.5)).
 However, formally verifying such certificates remains intractable due to the need to evaluate the network over a continuum of states.
 In the next section, we show how the system’s and certificate’s monotonicity enable safety verification and control synthesis from finite data.
-3.2 Monotone Barrier Certificates from Finite Data
+3.2. Monotone Barrier Certificates from Finite Data
 In this section, we exploit the system’s and certificates’ monotonicity to develop a data-driven framework for safety verification and control synthesis using only finite data, while ensuring formal correctness.
@@
 Given $\mathcal{X}_{0}\subseteq\mathcal{X}$ and $\mathcal{X}_{u}\subseteq\mathcal{X}$, we define index sets $\mathcal{I},\mathcal{U}$ as the minimal subsets of $\mathcal{P}$ such that
-$\mathcal{X}_{0}\subseteq\bigcup_{i\in\mathcal{I}}[\underline{x}^{i},\overline{x}^{i}],\qquad\mathcal{X}_{u}\subseteq\bigcup_{i\in\mathcal{U}}[\underline{x}^{i},\overline{x}^{i}].$ (6)
+(3.1) $\mathcal{X}_{0}\subseteq\bigcup_{i\in\mathcal{I}}[\underline{x}^{i},\overline{x}^{i}],\qquad\mathcal{X}_{u}\subseteq\bigcup_{i\in\mathcal{U}}[\underline{x}^{i},\overline{x}^{i}].$
 By construction, any $x\in\mathcal{X}_{0}$ satisfies $x\leq\overline{x}^{i}$ for some $i\in\mathcal{I}$, and any $x\in\mathcal{X}_{u}$ satisfies $\underline{x}^{i}\leq x$ for some $i\in\mathcal{U}$.
 For monotone systems with monotone barrier certificates, the points $\{\underline{x}^{i}\}_{i\in\mathcal{U}}$ and $\{\overline{x}^{i}\}_{i\in\mathcal{I}}$ act as worst-case representatives.
-Thus, the search for robust barrier certificates satisfying (2)-(4) can be conservatively reformulated as finding a monotone neural network $N:\mathcal{X}\to\mathbb{R}$ satisfying:
-$N(\overline{x}^{i})$ $\leq\gamma\quad$ $\forall i\in\mathcal{I},$ (7)
-$N(\underline{x}^{i})$ $>\eta\quad$ $\forall i\in\mathcal{U},$ (8)
-$N(\underline{x}^{i})$ $\leq\delta\quad\implies\quad\max_{j\in\mathcal{S}}N(f(\overline{x}^{i},\overline{u}^{j}))\leq\delta\quad$ $\forall i\in\mathcal{P}.$ (9)
-A monotone neural network $N:\mathcal{X}\to$ satisfying (7), (8), and (9) is called a monotone neural robust barrier certificate.
-Similarly, for safe control synthesis, the search can be reformulated as finding a monotone neural network $N:\mathcal{X}\to\mathbb{R}$ satisfying (7), (8), and
-$N(\underline{x}^{i})$ $\leq\delta\quad\implies\quad\min_{j\in\mathcal{S}}N(f(\overline{x}^{i},\underline{u}^{j}))\leq\delta\quad$ $\forall i\in\mathcal{P}.$ (10)
-A monotone neural network $N:\mathcal{X}\to$ that satisfies conditions (7), (8), and (9) is referred to as a monotone neural control barrier certificate.
-Conditions (7)–(10) reduce the verification of neural barrier certificates from the entire (continuous) state set to a finite set of carefully selected maximal and minimal points derived from the coverings of $\mathcal{X}$ and $U$.
+Thus, the search for robust barrier certificates satisfying (2.2)-(2.4) can be conservatively reformulated as finding a monotone neural network $N:\mathcal{X}\to\mathbb{R}$ satisfying:
+(3.2) $N(\overline{x}^{i})$ $\leq\gamma\quad$ $\forall i\in\mathcal{I},$
+(3.3) $N(\underline{x}^{i})$ $>\eta\quad$ $\forall i\in\mathcal{U},$
+(3.4) $N(\underline{x}^{i})$ $\leq\delta\quad\implies\quad\max_{j\in\mathcal{S}}N(f(\overline{x}^{i},\overline{u}^{j}))\leq\delta\quad$ $\forall i\in\mathcal{P}.$
+A monotone neural network $N:\mathcal{X}\to\mathbb{R}$ satisfying (3.2), (3.3), and (3.4) is called a monotone neural robust barrier certificate.
+Similarly, for safe control synthesis, the search can be reformulated as finding a monotone neural network $N:\mathcal{X}\to\mathbb{R}$ satisfying (3.2), (3.3), and
+(3.5) $N(\underline{x}^{i})$ $\leq\delta\quad\implies\quad\min_{j\in\mathcal{S}}N(f(\overline{x}^{i},\underline{u}^{j}))\leq\delta\quad$ $\forall i\in\mathcal{P}.$
+A monotone neural network $N:\mathcal{X}\to\mathbb{R}$ that satisfies conditions (3.2), (3.3), and (3.4) is referred to as a monotone neural control barrier certificate.
+Conditions (3.2)–(3.5) reduce the verification of neural barrier certificates from the entire (continuous) state set to a finite set of carefully selected maximal and minimal points derived from the coverings of $\mathcal{X}$ and $U$.
 The following result provides formal guarantees for safety verification and safe control synthesis using monotone neural barrier certificates.
-Theorem 4 (Monotone Neural Barrier Certificates).
+Theorem 3.2 (Monotone Neural Barrier Certificates).
 Consider a monotone dtDS $\mathfrak{S}=(\mathcal{X},\mathcal{X}_{0},U,f)$ with an unknown state transition map $f$ and a set of unsafe states $\mathcal{X}_{u}$.
-Let $\{[\underline{x}^{i},\overline{x}^{i}]\}_{i\in\mathcal{P}}$ be a family of hyper-rectangles that cover the state set $\mathcal{X}$, $\{[\underline{u}^{i},\overline{u}^{i}]\}_{i\in\mathcal{S}}$ be a family of hyper-rectangles that cover the input set $U$, and the index sets $\mathcal{I},\mathcal{U}$ be the minimal subsets of $\mathcal{P}$ satisfying (6).
+Let $\{[\underline{x}^{i},\overline{x}^{i}]\}_{i\in\mathcal{P}}$ be a family of hyper-rectangles that cover the state set $\mathcal{X}$, $\{[\underline{u}^{i},\overline{u}^{i}]\}_{i\in\mathcal{S}}$ be a family of hyper-rectangles that cover the input set $U$, and the index sets $\mathcal{I},\mathcal{U}$ be the minimal subsets of $\mathcal{P}$ satisfying (3.1).
 Then,
-  1. 1.
-    (Safety verification) If there exists a monotone neural network $N:\mathcal{X}\to$ satisfying (7), (8), and (9), with $\gamma\leq\delta\leq\eta$, then $\mathfrak{S}$ is safe with respect to $\mathcal{X}_{u}$, for every input $\{u_{t}\}_{t=0}^{\infty}\in U$;
-  2. 2.
-    (Safe control synthesis) If there exists a monotone neural network $N:\mathcal{X}\to$ satisfying (7), (8), and (10), with $\gamma\leq\delta\leq\eta$, then there exists a map $K:\mathcal{X}\to U$ such that $\mathfrak{S}$ with the feedback controller $\{u_{t}=K(x_{t})\}_{t=0}^{\infty}$ is safe with respect to $\mathcal{X}_{u}$.
-Proof. We refer to Appendix E for the proof.
-Theorem 4 highlights key advantages of our data-driven approach.
-It allows constructing safety certificates from arbitrary state space partitions, avoiding the curse of dimensionality in fine discretizations anand2023formally and the reliance on SMT solvers for neural network verification edwards2024fossil .
+  1. (1)
+    (Safety verification) If there exists a monotone neural network $N:\mathcal{X}\to\mathbb{R}$ satisfying (3.2), (3.3), and (3.4), with $\gamma\leq\delta\leq\eta$, then $\mathfrak{S}$ is safe with respect to $\mathcal{X}_{u}$, for every input $\{u_{t}\}_{t=0}^{\infty}\in U$;
+  2. (2)
+    (Safe control synthesis) If there exists a monotone neural network $N:\mathcal{X}\to\mathbb{R}$ satisfying (3.2), (3.3), and (3.5), with $\gamma\leq\delta\leq\eta$, then there exists a map $K:\mathcal{X}\to U$ such that $\mathfrak{S}$ with the feedback controller $\{u_{t}=K(x_{t})\}_{t=0}^{\infty}$ is safe with respect to $\mathcal{X}_{u}$.
+Proof. We refer to Appendix B for the proof.
+Theorem 3.2 highlights key advantages of our data-driven approach.
+It allows constructing safety certificates from arbitrary state space partitions, avoiding the curse of dimensionality in fine discretizations [3] and the reliance on SMT solvers for neural network verification [14].
 By leveraging system monotonicity, it scales to high-dimensional systems and does not require knowledge of Lipschitz constants for the dynamics or barrier function.
-3.3 Training Monotone Barrier Certificates.
+3.3. Training Monotone Barrier Certificates.
 In this section, we present an adaptive algorithm for training monotone neural robust (resp.
 control) barrier certificates.
-Let $\{[\underline{x}^{i},\overline{x}^{i}]\}_{i\in\mathcal{P}}$ and $\{[\underline{u}^{i},\overline{u}^{i}]\}_{i\in\mathcal{S}}$ be hyper-rectangular coverings of the state and input spaces, respectively, and $\mathcal{I},\mathcal{U}$ the index sets defined in (6).
+Let $\{[\underline{x}^{i},\overline{x}^{i}]\}_{i\in\mathcal{P}}$ and $\{[\underline{u}^{i},\overline{u}^{i}]\}_{i\in\mathcal{S}}$ be hyper-rectangular coverings of the state and input spaces, respectively, and $\mathcal{I},\mathcal{U}$ the index sets defined in (3.1).
 For safety verification, we train a monotone neural network $N:\mathcal{X}\to\mathbb{R}$ as a robust barrier certificate by minimizing the composite mean-squared-error loss $L_{\mathrm{total}}:=L_{1}+L_{2}+L_{3}$, where each term enforces a specific barrier certificate constraint:
-$L_{1}$ $:=\mathrm{MSE}(N(\overline{x}^{i}),\gamma),$ $\text{for all }i\in\mathcal{I},$ (11)
-$L_{2}$ $:=\mathrm{MSE}(N(\underline{x}^{i}),\eta),$ $\text{for all }i\in\mathcal{U},$ (12)
-$L_{3}$ $:=\mathrm{MSE}\left(\min_{j\in\mathcal{S}}N(f(\overline{x}^{i},\underline{u}^{j})),\delta\right),$ $\text{for all }i\in\mathcal{P}.$ (13)
+(3.6) $L_{1}$ $:=\mathrm{MSE}(N(\overline{x}^{i}),\gamma),$ $\text{for all }i\in\mathcal{I},$
+(3.7) $L_{2}$ $:=\mathrm{MSE}(N(\underline{x}^{i}),\eta),$ $\text{for all }i\in\mathcal{U},$
+(3.8) $L_{3}$ $:=\mathrm{MSE}\left(\min_{j\in\mathcal{S}}N(f(\overline{x}^{i},\underline{u}^{j})),\delta\right),$ $\text{for all }i\in\mathcal{P}.$
 Training uses gradient-based optimization on $L_{\mathrm{total}}$, with tunable parameters $\gamma,\delta,\eta\in\mathbb{R}_{>0}$ satisfying $\gamma<\delta<\eta$.
-The process stops once $N$ meets conditions (7), (8), and (9).
+The process stops once $N$ meets conditions (3.2), (3.3), and (3.4).
 For safe control synthesis, we jointly train two monotone neural networks: $N:\mathcal{X}\to\mathbb{R}$, which serves as the control barrier certificate, and $K:\mathcal{X}\to\mathcal{U}$, which represents the feedback policy.
-The objective for both of these networks is the composite mean-squared-error loss $\widehat{L}_{\mathrm{total}}:=L_{1}+L_{2}+L_{4}$, where $L_{1}$ and $L_{2}$ are defined in (11) and (12), and $L_{4}$ is given by
-$L_{4}$ $:=\mathrm{MSE}\left(N(f(\overline{x}^{i},K(\overline{x}^{i}))),\delta\right),$ $\forall i\in\mathcal{P}.$ (14)
+The objective for both of these networks is the composite mean-squared-error loss $\widehat{L}_{\mathrm{total}}:=L_{1}+L_{2}+L_{4}$, where $L_{1}$ and $L_{2}$ are defined in (3.6) and (3.7), and $L_{4}$ is given by
+(3.9) $L_{4}$ $:=\mathrm{MSE}\left(N(f(\overline{x}^{i},K(\overline{x}^{i}))),\delta\right),$ $\mbox{ for all }i\in\mathcal{P}.$
 Training is performed jointly on $N$ and $K$ using gradient-based optimization on $\widehat{L}_{\mathrm{total}}$, with tunable parameters $\gamma,\delta,\eta\in\mathbb{R}_{>0}$ such that $\gamma\leq\delta\leq\eta$.
-It stops when $N$ satisfies (7), (8), and the following condition:
+It stops when $N$ satisfies (3.2), (3.3), and the following condition:
 $N(\underline{x}^{i})\leq\delta\quad\Longrightarrow\quad N(f(\overline{x}^{i},K(\overline{x}^{i})))\leq\delta,\qquad\forall i\in\mathcal{P}.$
-$x_{1}$$x_{2}$
-(a) Initial Sets
-$x_{1}$$x_{2}$
-(b) Sampled Points
-$x_{1}$$x_{2}$
-(c) Refined Sets
-Figure 1: Illustration of hyper-rectangular cover refinement for a simple 2D system.
-The blue and red regions represent the initial and unsafe sets, respectively, while the purple region indicates where condition (10) must be satisfied.
-Two randomly selected sample points (black dots) are used to update the region where condition (10) is enforced.
-The condition is then verified at the maximal points of this region, shown as green dots.
-Iterative Refinement of the Hyper-rectangular Cover.
-In practice, the sample complexity for learning monotone neural barrier certificates depends on the granularity of the hyper-rectangular cover $\{[\underline{x}^{i},\overline{x}^{i}]\}_{i\in\mathcal{P}}$ of the state set $\mathcal{X}$.
-A fine hyper-rectangular cover for $\mathcal{X}$ can make enforcing  (10) across all cells computationally demanding.
-To address this, we propose an algorithm with linear sample complexity that starts with a coarse hyper-rectangular cover and refines it iteratively as needed to ensure  (10) always holds.
-Assume that $\mathcal{X}$ is a bounded hyper-rectangle $\mathcal{X}=[\underline{x}_{\min},\overline{x}_{\max}]$ which we consider as a hyper-rectangular cover of $\mathcal{X}$.
-We also suppose that there exists $\underline{x}_{u}\in\mathcal{X}$ such that $\mathcal{X}_{u}\subseteq[\underline{x}_{u},\overline{x}_{\max}]$.
-We define the set $T\subseteq\mathcal{X}$ by
-$T=[\underline{x}_{u},\overline{x}_{\max}],$ (15)
-and compute the set of all maximal points of $H=\mathcal{X}\setminus T$, denoted by $\{\overline{x}^{i}\}_{i\in\mathcal{P}}$ [^2] We refer to Appendix G for details about computing all maximal points of union of hyper-rectangles..
-In the first step, our algorithm refines the hyper-rectangular cover $\underline{x}_{\min},\overline{x}_{\max}]$ by replacing it with the family of hyper-rectangles $\{[\underline{x}_{\min},\overline{x}^{i}]\}_{i\in\mathcal{P}}$, and uses this family to define the loss function $L_{\mathrm{total}}$ for training a monotone neural network $N$.
-If $N$ fails to satisfy condition (10) after a set number of epochs, we sample $m$ intermediate states where the barrier certificate lies between the safe and unsafe thresholds and define the set $Q$ as follows:
-$Q:=m\mbox{ samples from }\left\{x\in\mathcal{X}\mid\gamma<N(x)<\eta\right\}.$
-Using the sampled points $x\in Q$, we update the set $T$ in (15) as follows:
-$T\leftarrow T\cup\bigcup_{x\in Q}[x,\overline{x}_{\max}],$
-Then we find all the maximal points of the set $\mathcal{X}\setminus T$ and update the family of hyper-rectangles $\{[\underline{x}_{\min},\overline{x}^{i}]\}_{i\in\mathcal{P}}$  [^3] We refer to Appendix G for details about computing all maximal points of union of hyper-rectangles..
-The barrier certificate is then retrained on the refined family of hyper-rectangles, where condition (10) is more likely to hold.
-The process repeats until all barrier conditions are met over the updated family of hyper-rectangles or a maximum number of iterations is reached.
-Figure 1 illustrates this iterative update of the family of hyper-rectangles in a 2D example.
-Algorithm 1 details the safe control synthesis, and Algorithm 2 covers safety verification.
-By leveraging monotonicity and carefully designing hyper-rectangular covers of the state space $\mathcal{X}$, our algorithm certifies the safety of unknown monotone systems with linear sample complexity.
-Algorithm 1 Monotone Neural Control Barrier Certificate
-Input: Simulator access to $f$, sets $\mathcal{X}_{0}$, $\mathcal{X}_{u}$, $\mathcal{X}=[\underline{x}_{\min},\overline{x}_{\max}]$, $e$ total number of epochs, $\hat{e}$ number of epochs before sampling, $m$ number of samples;
-Output: Neural networks $N$ (barrier certificate), $K$ (feedback controller);
-1:Over-approximate $\mathcal{X}_{0}$, $\mathcal{X}_{u}$, and $\mathcal{X}$ using unions of hyper-rectangles as in (6)
-2:Construct maximal initial points $\{\overline{x}^{i}\}_{i\in\mathcal{I}}$ and minimal unsafe points $\{\underline{x}^{i}\}_{i\in\mathcal{U}}$
-3:Construct the hyper-rectangular cover $T=[\underline{x}_{u},\overline{x}_{\max}]$
-4:Construct $H$ as the set of maximal points of $\mathcal{X}\setminus T$
-5:Initialize monotone neural networks NN, $K$ and set hyperparameters $\gamma<\delta=\eta$
-6:for $e$ number of epochs do
-7:  Minimize loss $\widehat{L}_{\mathrm{total}}=L_{1}+L_{2}+L_{4}$ as defined in (11), (12) and (14), respectively.
-8:  if Conditions (7)–(9) not satisfied for $\hat{e}$ epochs then
-9:   $Q\leftarrow$ Sample $m$ points from $\{x\in\mathcal{X}\mid\gamma<B(x)<\eta\}$
-10:   Update $T\leftarrow T\cup\bigcup_{x\in Q}[x,\overline{x}_{\max}]$
-11:   Update $H$ by getting maximal points of $\mathcal{X}\setminus T$
-12:   Set $\delta\leftarrow\min_{x\in H}N(x)$
-13:  end if
-14:  if all constraints are satisfied then
-15:   return trained networks NN, $K$
-16:  end if
-17:end for
-18:return Inconclusive (training failed to converge)
-Theorem 5 (Correctness and Sample Complexity of Algorithm 1).
-Consider a monotone (unknown) dtDS $\mathfrak{S}=(\mathcal{X},\mathcal{X}_{0},U,f)$ with $\mathcal{X}\subseteq\mathbb{R}^{n}$ and an upper closed unsafe set $\mathcal{X}_{u}\subseteq\mathcal{X}$.
-If Algorithm 1 returns a control barrier certificate $N:\mathcal{X}\to\mathbb{R}$ and a controller $K:\mathcal{X}\to U$, then the algorithm certifies the safety of $\mathfrak{S}$ under the feedback controller $\{u_{t}=K(x_{t})\}_{t=0}^{\infty}$ with respect to the unsafe set $\mathcal{X}_{u}$ using $\mathcal{O}(n)$ simulator queries.
-Proof. We refer to Appendix G for the proof.
-4 Experiments
-This section demonstrates the efficiency of our framework through three case studies, using an RTX 4090 GPU.
-Both the barrier certificate and feedback controller neural networks consist of $5$ hidden layers with $200$ neurons in each layer and ReLU activation functions.
-The monotone neural networks are trained using Algorithm 1 with $\hat{e}=1000$ and $e=20000$.
-While system models are mentioned, they are not used to encode barrier conditions.
-Table 1 compares training time and sample complexity of Algorithm 1 against state-of-the-art methods, and Table 2 shows scalability of Algorithm 1 for a traffic network from 2 to 100 states.
-Competing methods either time out after 12 hours or require exponentially many samples, while ours scales efficiently with linear sample growth.
-Edwards et al.
-(edwards2024fossil, ) rely on fixed random sampling and SMT-based verification (e.g., Z3 (de2008z3, )), which becomes intractable as dimensions increase.
-Table 1: Run time and sample complexity across different methods and case studies. All results are averaged over 30 random initializations. NA indicates failure to converge within 12 hours.
-Case Study (size) Edwards et al. edwards2024fossil Anand et al. anand2023formally Zhang et al. zhang2024seev Ours
-Running Time
-Traffic Control (50D) NA NA NA 20s
-Power Grid (13,659D) NA NA NA 250s
-Traffic Network (1000D) NA NA NA 135s
-Sample Complexity
-Traffic Control (50D) 500 $\Omega(2^{50})$ $\Omega(2^{50})$ 51
-Power Grid (13,659D) 500 $\Omega(2^{13000})$ $\Omega(2^{13000})$ 13,500
-Traffic Network (1000D) 500 $\Omega(2^{1000})$ $\Omega(2^{1000})$ 1001
-Table 2: Performance comparison on the Freeway Traffic Network as system dimension grows.
-Our method scales efficiently, while baselines either fail to converge or require exponential samples.
-Results are averaged over 30 runs.
-Dim Edwards et al. edwards2024fossil Anand et al. anand2023formally Zhang et al. zhang2024seev Ours
-Time Samples Time Samples Time Samples Time Samples
-2 30s 500 10m $10^{2}$ 12m $10^{2}$ 1s 3
-5 50s 500 45m $10^{5}$ 50m $10^{5}$ 2s 6
-10 2m 500 NA $\Omega(2^{10})$ NA $\Omega(2^{10})$ 4s 11
-25 NA 500 NA $\Omega(2^{25})$ NA $\Omega(2^{25})$ 10s 26
-50 NA 500 NA $\Omega(2^{50})$ NA $\Omega(2^{50})$ 20s 51
-75 NA 500 NA $\Omega(2^{75})$ NA $\Omega(2^{75})$ 45s 76
-100 NA 500 NA $\Omega(2^{100})$ NA $\Omega(2^{100})$ 90s 101
-Urban Traffic Control.
-Our first case study considers urban traffic control from sadraddini2018formal .
-The traffic network is modeled as a directed graph $(\mathcal{V},\mathcal{L})$, where each link $l\in\mathcal{L}$ supports one-way flow from tail node $\tau(l)\in\mathcal{V}\cup\{0\}$ to head node $\eta(l)\in\mathcal{V}$, with $\tau(l)=0$ representing external inflows.
-Our method synthesized a controller in 20s (averaged over 30 runs).
-Details are in Appendix H.
-European Power Grid Verification.
-Our second benchmark is the 13659-bus European grid from the pegase project josz2016ac , with lossless swing dynamics discretized as
-$\theta_{k+1}=\theta_{k}+\tau\bigl{(}P-BA\sin(B^{\!\top}\theta_{k})\bigr{)}:=f(\theta_{k},P),$ (16)
-where $\tau$ is the step size, $\theta\in\mathbb{R}^{n}$ are bus angles, $P\in^{n}$ are power generations/demands at buses, $B$ is the incidence matrix of the interconnection power grid, and $A=\operatorname{diag}(V_{i}V_{j}Y_{ij})$, where $Y_{ij}$ is the line admittances and $V_{i}$ is the voltage amplitude at bus $i$.
-This power system is a monotone dtDS $\mathfrak{S}=(\mathcal{X},\mathcal{X}_{0},U,f)$ with the state set $\mathcal{X}=[\frac{\pi}{2},\frac{\pi}{2}]^{n}$, the initial set $\mathcal{X}_{0}=[0,0.1]^{n}$, the input set $U=\{P_{\mathrm{nom}}\}$, and the unsafe set $\mathcal{X}_{u}=[\frac{\pi}{4},\frac{\pi}{2}]^{n}$ with $n=13659$, and $\tau=0.1$.
-Using the Matpower black-box model zimmerman1997matpower , our method certifies safety in 250s on average over 30 runs.
-Freeway Traffic Verification.
-Finally, we consider a Daganzo cell transmission model with demand and supply constraints, as shown in Figure 4.
-The traffic network is modeled as a monotone system $\mathfrak{S}=(\mathcal{X},\mathcal{X}_{0},U,f)$ FB-CTDS with the state set $\mathcal{X}=[0,50]^{n}$, the initial set $\mathcal{X}_{0}=[0,5]\times[0,18]\times[0,5]^{n-2}$, the input $U=[0,9]$ and the unsafe set is $\mathcal{X}_{u}=[30,50]^{n}$.
-Our method finds a monotone neural robust barrier certificate in 135s for a 1000-dimensional network (averaged over 30 runs).
-See Appendix I for details.
-5 Conclusions
+4. Conclusions
 This work advances neurosymbolic verification by proposing a scalable framework for synthesizing safety controllers in high-dimensional monotone systems using monotone neural barrier certificates.
-Leveraging the systems’ order-preserving structure, our approach reduces verification to local boundary checks, achieving potentially linear sample complexity without explicit models or conservative assumptions.
-Experiments on large-scale benchmarks demonstrate its practicality and scalability.
-Future work may extend the framework to systems with partial or mixed monotonicity and explore integration with reinforcement learning for adaptive control.
-NeurIPS Paper Checklist
-  1. 1.
-    Claims
-  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
-  3. Answer: [Yes]
-  4. Justification: The claims match theoretical and experimental results.
-  5. Guidelines:
-    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
-    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations.
-A No or NA answer to this question will not be perceived well by the reviewers.
-    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
-    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
-  6. 2.
-    Limitations
-  7. Question: Does the paper discuss the limitations of the work performed by the authors?
-  8. Answer: [Yes]
-  9. Justification: We have explicitly mentioned that this work only supports monotone systems.
-  10. Guidelines:
-    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
-    • The authors are encouraged to create a separate "Limitations" section in their paper.
-    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally).
-The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
-    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs.
-In general, empirical results often depend on implicit assumptions, which should be articulated.
-    • The authors should reflect on the factors that influence the performance of the approach.
-For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting.
-Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
-    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
-    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
-    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper.
-The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community.
-Reviewers will be specifically instructed to not penalize honesty concerning limitations.
-  11. 3.
-    Theory assumptions and proofs
-  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
-  13. Answer: [Yes]
-  14. Justification: Every theorem is backed by its proof (if novel) or referenced.
-  15. Guidelines:
-    • The answer NA means that the paper does not include theoretical results.
-    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
-    • All assumptions should be clearly stated or referenced in the statement of any theorems.
-    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
-    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
-    • Theorems and Lemmas that the proof relies upon should be properly referenced.
-  16. 4.
-    Experimental result reproducibility
-  17.
-Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
-  18. Answer: [Yes]
-  19. Justification: We have included architecture of neural networks, dynamical systems, and hyper-parameters. Code is also available.
-  20. Guidelines:
-    • The answer NA means that the paper does not include experiments.
-    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
-    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
-    • Depending on the contribution, reproducibility can be accomplished in various ways.
-For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model.
-In general.
-releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
-    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution.
-For example
-      1. (a)
-        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
-      2. (b)
-        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
-      3. (c)
-        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
-      4. (d)
-        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility.
-In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
-  21. 5.
-    Open access to data and code
-  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
-  23. Answer: [Yes]
-  24. Justification: Code is available.
-  25. Guidelines:
-    • The answer NA means that paper does not include experiments requiring code.
-    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
-    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer.
-Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
-    • The instructions should contain the exact command and environment needed to run to reproduce the results.
-See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
-    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
-    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines.
-If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
-    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
-    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
-  26. 6.
-    Experimental setting/details
-  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
-  28. Answer: [Yes]
-  29. Justification: Every detail is mentioned in experiments’ section.
-  30. Guidelines:
-    • The answer NA means that the paper does not include experiments.
-    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
-    • The full details can be provided either with the code, in appendix, or as supplemental material.
-  31. 7.
-    Experiment statistical significance
-  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
-  33. Answer: [No]
-  34. Justification: We have formal proof, which is deterministic.
-  35. Guidelines:
-    • The answer NA means that the paper does not include experiments.
-    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
-    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
-    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
-    • The assumptions made should be given (e.g., Normally distributed errors).
-    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
-    • It is OK to report 1-sigma error bars, but one should state it.
-The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
-    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
-    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
-  36. 8.
-    Experiments compute resources
-  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
-  38. Answer: [Yes]
-  39. Justification: We trained on a single RTX 4090, as mentioned in the paper.
-  40. Guidelines:
-    • The answer NA means that the paper does not include experiments.
-    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
-    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
-    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
-  41. 9.
-    Code of ethics
-  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
-  43. Answer: [Yes]
-  44. Justification: We have faithfully followed Code of Ethics.
-  45. Guidelines:
-    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
-    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
-    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
-  46. 10.
-    Broader impacts
-  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
-  48. Answer: [Yes]
-  49. Justification: We have disclosed our broader impact. Our work paves the way for AI agents to be implemented in real-life scenarios with safety guarantees.
-  50. Guidelines:
-    • The answer NA means that there is no societal impact of the work performed.
-    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
-    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
-    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments.
-However, if there is a direct path to any negative applications, the authors should point it out.
-For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation.
-On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
-    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
-    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
-  51. 11.
-    Safeguards
-  52.
-Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
-  53. Answer: [N/A]
-  54. Justification: Our method ensures safety.
-  55. Guidelines:
-    • The answer NA means that the paper poses no such risks.
-    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
-    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
-    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
-  56. 12.
-    Licenses for existing assets
-  57.
-Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
-  58. Answer: [Yes]
-  59. Justification: We have followed the guidelines.
-  60. Guidelines:
-    • The answer NA means that the paper does not use existing assets.
-    • The authors should cite the original paper that produced the code package or dataset.
-    • The authors should state which version of the asset is used and, if possible, include a URL.
-    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
-    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
-    • If assets are released, the license, copyright information, and terms of use in the package should be provided.
-For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets.
-Their licensing guide can help determine the license of a dataset.
-    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
-    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
-  61. 13.
-    New assets
-  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
-  63. Answer: [Yes]
-  64. Justification: We have included every detail about our models and training.
-  65. Guidelines:
-    • The answer NA means that the paper does not release new assets.
-    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
-    • The paper should discuss whether and how consent was obtained from people whose asset is used.
-    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
-  66. 14.
-    Crowdsourcing and research with human subjects
-  67.
-Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
-  68. Answer: [N/A]
-  69. Justification: Paper does not involve crowdsourcing nor research with human subjects.
-  70. Guidelines:
-    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
-    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
-    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
-  71. 15.
-    Institutional review board (IRB) approvals or equivalent for research with human subjects
-  72.
-Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
-  73. Answer: [N/A]
-  74. Justification: The paper does not involve crowdsourcing nor research with human subjects.
-  75. Guidelines:
-    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
-    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research.
-If you obtained IRB approval, you should clearly state this in the paper.
-    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
-    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
-  76. 16.
-    Declaration of LLM usage
-  77.
-Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research?
-Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
-  78. Answer: [N/A]
-  79. Justification: LLMs were used only for editing.
-  80. Guidelines:
-    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
-    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.
-Appendix A Related Work
-Barrier Certificates.
-[38] introduced barrier certificates (BCs) as functions whose level sets over-approximate system reachable sets.
-Control barrier certificates (CBCs) later emerged as a key tool for synthesizing safe controllers [3, 49, 24, 11, 15], typically using Sum-of-Squares (SOS) optimization [39, 57, 44].
-Neural Barrier Certificates.
-[55] first introduced neural network–based barrier certificates, using a shallow network and MILP for formal verification.
-Later, [35] employed SMT solvers to find counterexamples for iterative refinement.
-Neural barrier certificates have since been extended to hybrid systems [56] and stochastic dynamics [32].
-Neural Control Barrier Certificates.
-Neural control barrier certificates (NCBCs) extend these ideas to safe controller synthesis.
-Recent work [41, 29, 18, 17, 31, 30] learns both barrier functions and control policies using neural networks.
-Verification typically relies on SMT solvers [1, 55, 20], reachable set methods [48], polynomial approximations [45], Lipschitz bounds [5], or ReLU verification [26].
-Recently, [54, 53] an exact NCBC verification algorithm was proposed that synthesizes a controller if the barrier is valid.
-Most of these methods, however, assume full system knowledge (except [5]) and rely heavily on SMT solvers, which scale poorly to deep networks due to their exponential complexity.
-Monotone dynamics.
-Monotone systems are dynamical systems whose trajectories preserve a partial order, exhibiting regular behaviors common in applications like traffic networks [14] and biological systems [47].
-While monotonicity has been shown to reduce sample complexity in Lyapunov analysis [40], its role in constructing barrier certificates remains less explored.
-Appendix B Proof of Theorem 1
-Appendix C Proof of Theorem 2
-Appendix D Proof of Theorem 3
-Before proving Theorem 3 part 1, we introduce several notations and intermediate results that will be used throughout the proof.
-Consider a monotone dtDS $\mathfrak{S}=(\mathcal{X},\mathcal{X}_{0},U,f)$.
-The reachable set of $\mathfrak{S}$, denoted by $\Omega\subseteq\mathcal{X}$, is defined as follows:
-$\Omega$ $=\bigcup_{T\in\mathbb{Z}_{\geq 0}}\Omega_{T},$ (17)
-[…]
-We first state the following intermediate result about the reachable set $\Omega$.
-Lemma 1.
-Consider a monotone dtDS $\mathfrak{S}=(\mathcal{X},\mathcal{X}_{0},U,f)$ with the reachable set $\Omega$ defined in (17). Then, for every $x\in\Omega$ and every $u\in U$, we have $f(x,u)\in\Omega$.
-Proof.
-Let $x\in\Omega$ and $u\in U$.
-Then there exists $T\in\mathbb{Z}_{\geq 0}$ such that $x\in\Omega_{T}$.
-This means that there exists $\{x_{t}\}_{t=0}^{\infty}$ a trajectory of $\mathfrak{S}$ with the input $\{u_{t}\}_{t=0}^{\infty}$ starting from $x_{0}\in\mathcal{X}_{0}$ such that $x=x_{T}$.
-Now consider the input $\{w_{t}\}_{t=0}^{\infty}$ defined by
-$w_{t}=\begin{cases}u_{t}&t\leq T,\\ u&t>T.\end{cases}$ (18)
-Let $\{y_{t}\}_{t=0}^{\infty}$ be the trajectory of $\mathfrak{S}$ with the input $\{w_{t}\}_{t=0}^{\infty}$ starting from $x_{0}$.
-By definition of the input $\{w_{t}\}_{t=0}^{\infty}$ in (18), it is clear that $y_{t}=x_{t}$, for every $t\leq T$.
-Moreover, for $t=T+1$, we have
-$y_{T+1}=f(y_{T},w_{T})=f(y_{T},u)=f(x_{T},u)=f(x,u),$
-where the second equality holds by definition of the input $\{w_{t}\}_{t=0}^{\infty}$ in (18), the third equality holds using the fact that $x_{t}=y_{t}$, for every $t\leq T$, and the fourth equality holds because $x=x_{T}$.
-Therefore, by definition of $\Omega_{T+1}$, we have $f(x,u)\in\Omega_{T+1}$ and thus $f(x,u)\in\Omega$.
-∎
-Motivated by the definition of the reachable set of $\mathfrak{S}$ in (17) we define the set $\widehat{\Omega}\subseteq\mathcal{X}$ as the smallest lower closed set in $\mathcal{X}$ containing $\Omega$.
-More specifically, $\widehat{\Omega}$ is defined as:
-$\widehat{\Omega}=\{x\in\mathcal{X}\;|\;x\leq y,\mbox{ for some }y\in\Omega\}.$ (19)
-Given $\gamma,\eta\in_{\geq 0}$ such that $\gamma\leq\eta$, we use the lower closed set $\widehat{\Omega}$ to define the function $B:\mathcal{X}\to$ as follows:
-$B(x)=\begin{cases}\gamma&x\in\widehat{\Omega}\\ \eta+1&x\not\in\widehat{\Omega}\end{cases}$ (20)
-The next lemma proves a useful property of the function $B$.
-Lemma 2.
-Consider a monotone dtDS $\mathfrak{S}=(\mathcal{X},\mathcal{X}_{0},U,f)$ with the lower closed set $\widehat{\Omega}$ as defined in (19).
-Then the function $B:\mathcal{X}\to$ defined in (20) is monotone.
-Proof.
-Let $x,y\in\mathcal{X}$ be such that $x\leq y$. We consider two cases i) $y\in\widehat{\Omega}$ and ii) $y\not\in\widehat{\Omega}$.
-  1. i)
-    In this case, by definition of $\widehat{\Omega}$ in (19), there exists a point $z\in\Omega$ such that $y\leq z$.
-Moreover, we have $x\leq y\leq z$.
-Therefore, we have $x\in\widehat{\Omega}$ and $B(x)=\gamma=B(y)$.
-  2. ii)
-    In this case, we have $B(y)=\eta+1$. Since we know that $\gamma\leq\eta<\eta+1$, we get $B(x)\leq\eta+1=B(y)$.
-Thus, in both cases i) and ii), we have $B(x)\leq B(y)$. This implies that $B$ is a monotone map. ∎
-Now we go back to the proof of Theorem 3 part 1. We divide the proof into if part and only if part.
-Proof of if part.
-Suppose that $B:\mathcal{X}\to$ is a monotone function satisfying conditions (2),(3), and (4).
-By definition 2, $B:\mathcal{X}\to$ is a robust barrier certificate for $\mathfrak{S}$ with respect to $\mathcal{X}_{u}$.
-Therefore, by Theorem 1, $\mathfrak{S}$ is safe with respect to $\mathcal{X}_{u}$, for every input $\{u_{t}\}_{t=0}^{\infty}$.
-Proof of only if part.
-Now assume that the monotone dtDS $\mathfrak{S}$ is safe with respect to upper closed unsafe set $\mathcal{X}_{u}$, for every input $\{u_{t}\}_{t=0}^{\infty}\in U$.
-Let $\widehat{\Omega}$ be the lower closed set defined in (19) and $B:\mathcal{X}\to$ be the function defined in (20).
-Our goal is to show that $B:\mathcal{X}\to$ defined in (20) is a monotone robust barrier certificate for $\mathfrak{S}$ satisfying conditions (2),(3), and (4).
-First, by Lemma 2, the function $B:\mathcal{X}\to$ is monotone.
-  1. 1.
-    ($B$ satisfies condition (2)) Suppose that $x\in\mathcal{X}_{0}$.
-By the definition of $\Omega_{0}$, this implies that $x\in\Omega_{0}$.
-As a result, we have $x\in\Omega\subseteq\widehat{\Omega}$.
-Using the definition of $B$ in (20), we get $B(x)=\gamma$.
-This shows that $B$ satisfies condition (2).
-  2. 2.
-    ($B$ satisfies condition (3)) Suppose that $x\in\mathcal{X}_{u}$.
-Since $\mathfrak{S}$ is safe with respect to $\mathcal{X}_{u}$ for any input $\{u_{t}\}_{t=0}^{\infty}\in U$, we get that $x\not\in\Omega$.
-Note that $\widehat{\Omega}$ is the smallest lower closed set containing $\Omega$ and $\mathcal{X}_{u}$ is an upper closed set.
-As a result, $x\not\in\widehat{\Omega}$.
-By definition of $B$ in (20), we have $B(x)=\eta+1>\eta$.
-This shows that $B$ satisfies condition (3).
-  3. 3.
-    ($B$ satisfies condition (4)) Let $x\in\mathcal{X}$ be such that $B(x)\leq\delta$.
-Note that $\delta\leq\eta<\eta+1$.
-By definition of the function $B$ in (20), this implies that $x\in\widehat{\Omega}$.
-By definition of the set $\widehat{\Omega}$ in (19), there exists $y\in\Omega$ such that $x\leq y$.
-Since the state transition map $f$ is monotone, for every $u\in U$, we have $f(x,u)\leq f(y,u)$.
-Note that $y\in\Omega$ and $u\in U$.
-Therefore, using Lemma 1, we can show that $f(y,u)\in\Omega$.
-Using the definition of the lower closed set $\widehat{\Omega}$ in (19), this implies that $f(x,u)\in\widehat{\Omega}$.
-As a result, one can use the definition of $B$ in (20) to show that $B(f(x,u))=\gamma$.
-Since this equality holds, for every $u\in U$, we get
-    $\sup_{u\in U}B(f(x,u))=\gamma\leq\delta.$
-    This shows that $B$ satisfies condition (4).
-This completes the proof of Theorem 3 part 1.
-Now we focus on the proof of Theorem 3 part 2.
-Before we state the proof, we introduce several notations and intermediate results that will be used throughout the proof.
-Consider a monotone dtDS $\mathfrak{S}=(\mathcal{X},\mathcal{X}_{0},U,f)$ with a map $K:\mathcal{X}\to U$.
-The reachable set of $\mathfrak{S}$ with the feedback controller $u=K(x)$, which is denoted by $\Omega^{K}\subseteq\mathcal{X}$, is defined by
-$\Omega^{K}$ $=\bigcup_{T\in\mathbb{Z}_{\geq 0}}\Omega^{K}_{T},$ (21)
-[…]
-Motivated by the definition of the reachable set of $\mathfrak{S}$ in (17) we define the set $\widehat{\Omega}^{K}\subseteq\mathcal{X}$ as the smallest lower closed set in $\mathcal{X}$ containing $\Omega^{K}$.
-More specifically, $\widehat{\Omega}^{K}$ is defined as:
-$\widehat{\Omega}^{K}=\{x\in\mathcal{X}\;|\;x\leq y,\mbox{ for some }y\in\Omega^{K}\}.$ (22)
-Given $\gamma,\eta\in_{\geq 0}$ such that $\gamma\leq\eta$, we use the lower closed set $\widehat{\Omega}^{K}$ to define the function $B^{K}:\mathcal{X}\to$ as follows:
-$B^{K}(x)=\begin{cases}\gamma&x\in\widehat{\Omega}^{K}\\ \eta+1&x\not\in\widehat{\Omega}^{K}\end{cases}$ (23)
-The next lemma proves a useful property of the function $B^{K}$.
-Lemma 3.
-Consider a monotone dtDS $\mathfrak{S}=(\mathcal{X},\mathcal{X}_{0},U,f)$ with the lower closed set $\widehat{\Omega}$ as defined in (19).
-Then the function $B^{K}:\mathcal{X}\to$ defined in (23) is monotone.
-Proof.
-Let $x,y\in\mathcal{X}$ be such that $x\leq y$. We consider two cases: i) $y\in\widehat{\Omega}^{K}$, and ii) $y\not\in\widehat{\Omega}^{K}$.
-  1. 1.
-    In this case, there exists $z\in\Omega^{K}$ such that $y\leq z$.
-Note that we have $x\leq y\leq z$.
-This implies that $x\in\widehat{\Omega}^{K}$.
-As a result, by definition of $B^{K}$ in (23).
-we have $B^{K}(x)=B^{K}(y)=\gamma$.
-  2. 2.
-    In this case, by definition of $B^{K}$ in (23), we have $B^{K}(y)=\eta$. This implies that $B^{K}(y)=\eta+1$, since Note that $\gamma\leq\eta<\eta+1$. This means that $B(x)\leq\eta+1=B(y)$.
-In both cases i) and ii), we have $B^{K}(x)\leq B^{K}(y)$. This implies that $B^{K}$ is a monotone map. ∎
-Now we get back to Theorem 3 part 2. We divide the proof into if part and only if part.
-Proof of if part.
-Suppose that $B:\mathcal{X}\to$ is a monotone barrier certificate satisfying (2),(3), and (5).
-Then, by Theorem 2, there exists a map $K:\mathcal{X}\to$ such that the system $\mathfrak{S}$ with the feedback controller $\{u_{t}=K(x_{t})\}_{t=0}^{\infty}$ is safe with respect to $\mathcal{X}_{u}$.
-Proof of only if part.
-Suppose that there exists a map $K:\mathcal{X}\to$ such that the system $\mathfrak{S}$ with the feedback controller $\{u_{t}=K(x_{t})\}_{t=0}^{\infty}$ is safe with respect to $\mathcal{X}_{u}$.
-Let the lower closed set $\widehat{\Omega}^{K}$ be defined in (19) and the function $B^{K}:\mathcal{X}\to$ be defined as in 23.
-Our goal is to show that $B^{K}:\mathcal{X}\to$ is a monotone control barrier certificate for $\mathfrak{S}$ satisfying conditions (2),(3), and (5).
-By Lemma (3), it is clear that $B^{K}$ is monotone.
-  1. 1.
-    ($B^{K}$ satisfies condition (2)) Suppose that $x\in\mathcal{X}_{0}$.
-By the definition of $\Omega^{K}_{0}$, this implies that $x\in\Omega^{K}_{0}$.
-As a result, we have $x\in\Omega^{K}\subseteq\widehat{\Omega}^{K}$.
-Using the definition of $B^{K}$ in (23), we get $B^{K}(x)=\gamma$.
-This shows that $B^{K}$ satisfies condition (2).
-  2. 2.
-    ($B^{K}$ satisfies condition (3)) Suppose that $x\in\mathcal{X}_{u}$.
-Since $\mathfrak{S}$ with the feedback controller $u=K(x)$ is safe with respect to $\mathcal{X}_{u}$, we get that $x\not\in\Omega^{K}$.
-Note that $\widehat{\Omega}^{K}$ is the smallest lower closed set containing $\Omega^{K}$ and $\mathcal{X}_{u}$ is an upper closed set.
-As a result, $x\not\in\widehat{\Omega}^{K}$.
-By definition of $B^{K}$ in (23), we have $B^{K}(x)=\eta+1>\eta$.
-This shows that $B^{K}$ satisfies condition (3).
-  3. 3.
-    ($B^{K}$ satisfies condition (5)) Let $x\in\mathcal{X}$ be such that $B^{K}(x)\leq\delta$.
-Note that $\delta\leq\eta<\eta+1$.
-By definition of the function $B^{K}$ in (23), this implies that $x\in\widehat{\Omega}^{K}$.
-By definition of the set $\widehat{\Omega}^{K}$ in (22), there exists $y\in\Omega^{K}$ such that $x\leq y$.
-Since the state transition map $f$ is monotone, for every $u\in U$, we have $f(x,K(y))\leq f(y,K(y))$.
-Note that $y\in\Omega^{K}$.
-Therefore, by definition of $\Omega^{K}$ in (21), we have $f(y,K(y))\in\Omega^{K}$.
-Using the definition of the lower closed set $\widehat{\Omega}^{K}$ in (19), this implies that $f(x,K(y))\in\widehat{\Omega}$.
-As a result, we have
-    $\inf_{u\in U}B^{K}(f(x,u))\leq B^{K}(f(x,K(y)))=\gamma\leq\delta,$
-    where the first inequality holds by monotonicity of $f$ and $B^{K}$ and the second equality holds by definition of $B^{K}$ in (20). This shows that $B^{K}$ satisfies condition (5).
-This completes the proof of Theorem 3 part 2.
-Appendix E Proof of Theorem 4
-Regarding part 1, our goal is to show that the monotone neural network $N:\mathcal{X}\to$ satisfying conditions (7), (8), and (9) is a robust barrier certificate for $\mathfrak{S}$ with respect to the unsafe set $\mathcal{X}_{u}$.
-More specifically, our goal is to show that the monotone neural network $N:\mathcal{X}\to$ satisfies conditions (2), (3), and (4).
-  1. 1.
-    (NN satisfies condition (2)) Let $x\in\mathcal{X}_{0}$.
+Leveraging the systems’ order-preserving structure, our approach reduces verification of monotone systems to local boundary checks.
+Appendix A Proof of Theorem 3.1
+Suppose that $B:\mathcal{X}\to\mathbb{R}$ is a monotone function satisfying conditions (2.2),(2.3), and (2.4).
+By definition 2.2, $B:\mathcal{X}\to\mathbb{R}$ is a robust barrier certificate for $\mathfrak{S}$ with respect to $\mathcal{X}_{u}$.
+Therefore, by Theorem 2.3, $\mathfrak{S}$ is safe with respect to $\mathcal{X}_{u}$, for every input $\{u_{t}\}_{t=0}^{\infty}$.
+Now we focus on the proof of Theorem 3.1 part 2.
+Suppose that $B:\mathcal{X}\to\mathbb{R}$ is a monotone barrier certificate satisfying (2.2),(2.3), and (2.5).
+Then, by Theorem 2.5, there exists a map $K:\mathcal{X}\to\mathbb{R}$ such that the system $\mathfrak{S}$ with the feedback controller $\{u_{t}=K(x_{t})\}_{t=0}^{\infty}$ is safe with respect to $\mathcal{X}_{u}$.
+Appendix B Proof of Theorem 3.2
+Regarding part 1, our goal is to show that the monotone neural network $N:\mathcal{X}\to\mathbb{R}$ satisfying conditions (3.2), (3.3), and (3.4) is a robust barrier certificate for $\mathfrak{S}$ with respect to the unsafe set $\mathcal{X}_{u}$.
+More specifically, our goal is to show that the monotone neural network $N:\mathcal{X}\to\mathbb{R}$ satisfies conditions (2.2), (2.3), and (2.4).
+  1. (1)
+    (NN satisfies condition (2.2)) Let $x\in\mathcal{X}_{0}$.
 Since $\mathcal{X}_{0}\subseteq\bigcup_{i\in\mathcal{I}}[\underline{x}^{i},\overline{x}^{i}]$, there exists $k\in\mathcal{I}$ such that $x\leq\overline{x}^{k}$.
-Since $N:\mathcal{X}\to$ satisfies condition (7), we have $N(\overline{x}^{k})\leq\gamma$.
+Since $N:\mathcal{X}\to\mathbb{R}$ satisfies condition (3.2), we have $N(\overline{x}^{k})\leq\gamma$.
 On the other hand, NN is a monotone neural network, thus $N(x)\leq N(\overline{x}^{k})\leq\gamma$.
-This shows that the monotone neural network $N$ satisfies condition (2).
-  2. 2.
-    (NN satisfies condition (3)) Let $x\in\mathcal{X}_{u}$.
+This shows that the monotone neural network $N$ satisfies condition (2.2).
+  2. (2)
+    (NN satisfies condition (2.3)) Let $x\in\mathcal{X}_{u}$.
 Since $\mathcal{X}_{u}\subseteq\bigcup_{i\in\mathcal{U}}[\underline{x}^{i},\overline{x}^{i}]$, there exists $j\in\mathcal{U}$ such that $\underline{x}^{j}\leq x$.
-Since $N$ satisfies the condition (8), we have $\eta\leq N(\underline{x}^{j})$.
+Since $N$ satisfies the condition (3.3), we have $\eta\leq N(\underline{x}^{j})$.
 Using the monotonicity of the neural network NN, we get $\eta\leq N(\underline{x}^{j})\leq N(x)$.
-This means that the monotone neural network $N$ satisfies condition (3).
-  3. 3.
-    (NN satisfies condition (4)) Let $x\in\mathcal{X}$ be such that $N(x)\leq\delta$.
+This means that the monotone neural network $N$ satisfies condition (2.3).
+  3. (3)
+    (NN satisfies condition (2.4)) Let $x\in\mathcal{X}$ be such that $N(x)\leq\delta$.
 Since the family of hyper-rectangles $\{[\underline{x}^{i},\overline{x}^{i}]\}_{i\in\mathcal{P}}$ covers the state set $\mathcal{X}$, there exists $k\in\mathcal{P}$ such that $x\in[\underline{x}^{k},\overline{x}^{k}]$.
 Using the monotonicity of $N$, we get $N(\underline{x}^{k})\leq N(x)\leq\delta$.
-Since $N$ satisfies condition (9), we have $\max_{j\in\mathcal{S}}N(f(\overline{x}^{k},\overline{u}^{j}))\leq\delta$.
+Since $N$ satisfies condition (3.4), we have $\max_{j\in\mathcal{S}}N(f(\overline{x}^{k},\overline{u}^{j}))\leq\delta$.
 As a result, for every $u\in U$, we have
@@
     where the first inequality holds by the monotonicity of $N$ and $f$ and the fact that $x\leq\overline{x}^{k}$, the second inequality holds by monotonicity of $N$ and $f$ and the fact that $u\leq\max_{j\in\mathcal{S}}\overline{u}^{j}$.
-This means that $N$ satisfies condition (4).
-Therefore, the monotone neural network $N:\mathcal{X}\to$ satisfies conditions (2), (3), and (4). The result then follows using Theorem (1). This completes the proof of part 1.
-Regarding part 2, our goal is to show that the monotone neural network $N:\mathcal{X}\to$ satisfying conditions (7), (8), and (10) is a control barrier certificate for $\mathfrak{S}$ with respect to $\mathcal{X}_{u}$.
-More specifically, our goal is to show that $N$ satisfies conditions (2), (3), and (5).
-  1. 1.
-    (NN satisfies condition (2)) Let $x\in\mathcal{X}_{0}$.
+This means that $N$ satisfies condition (2.4).
+Therefore, the monotone neural network $N:\mathcal{X}\to\mathbb{R}$ satisfies conditions (2.2), (2.3), and (2.4). The result then follows using Theorem (2.3). This completes the proof of part 1.
+Regarding part 2, our goal is to show that the monotone neural network $N:\mathcal{X}\to\mathbb{R}$ satisfying conditions (3.2), (3.3), and (3.5) is a control barrier certificate for $\mathfrak{S}$ with respect to $\mathcal{X}_{u}$.
+More specifically, our goal is to show that $N$ satisfies conditions (2.2), (2.3), and (2.5).
+  1. (1)
+    (NN satisfies condition (2.2)) Let $x\in\mathcal{X}_{0}$.
 Since $\mathcal{X}_{0}\subseteq\bigcup_{i\in\mathcal{I}}[\underline{x}^{i},\overline{x}^{i}]$, there exists $k\in\mathcal{I}$ such that $x\leq\overline{x}^{k}$.
-Since $N:\mathcal{X}\to$ satisfies condition (7), we have $N(\overline{x}^{k})\leq\gamma$.
+Since $N:\mathcal{X}\to\mathbb{R}$ satisfies condition (3.2), we have $N(\overline{x}^{k})\leq\gamma$.
 On the other hand, NN is a monotone neural network, thus $N(x)\leq N(\overline{x}^{k})\leq\gamma$.
-This shows that the monotone neural network $N$ satisfies condition (2).
-  2. 2.
-    (NN satisfies condition (3)) Let $x\in\mathcal{X}_{u}$.
+This shows that the monotone neural network $N$ satisfies condition (2.2).
+  2. (2)
+    (NN satisfies condition (2.3)) Let $x\in\mathcal{X}_{u}$.
 Since $\mathcal{X}_{u}\subseteq\bigcup_{i\in\mathcal{U}}[\underline{x}^{i},\overline{x}^{i}]$, there exists $j\in\mathcal{U}$ such that $\underline{x}^{j}\leq x$.
-Since $N$ satisfies the condition (8), we have $\eta\leq N(\underline{x}^{j})$.
+Since $N$ satisfies the condition (3.3), we have $\eta\leq N(\underline{x}^{j})$.
 Using the monotonicity of the neural network NN, we get $\eta\leq N(\underline{x}^{j})\leq N(x)$.
-This means that the monotone neural network $N$ satisfies condition (3).
-  3. 3.
-    (NN satisfies condition (5)) Let $x\in\mathcal{X}$ be such that $N(x)\leq\delta$.
+This means that the monotone neural network $N$ satisfies condition (2.3).
+  3. (3)
+    (NN satisfies condition (2.5)) Let $x\in\mathcal{X}$ be such that $N(x)\leq\delta$.
 Since the family of hyper-rectangles $\{[\underline{x}^{i},\overline{x}^{i}]\}_{i\in\mathcal{P}}$ covers the state set $\mathcal{X}$, there exists $k\in\mathcal{P}$ such that $x\in[\underline{x}^{k},\overline{x}^{k}]$.
 Using the monotonicity of $N$, we get $N(\underline{x}^{k})\leq N(x)\leq\delta$.
-Since $N$ satisfies condition (10), we have $\min_{j\in\mathcal{S}}N(f(\overline{x}^{k},\overline{u}^{j}))\leq\delta$.
+Since $N$ satisfies condition (3.5), we have $\min_{j\in\mathcal{S}}N(f(\overline{x}^{k},\overline{u}^{j}))\leq\delta$.
 Since both $N$ and $f$ are monotone functions, for every $u\in U$, we have $N(f(x,u))\leq N(f(\overline{x}^{k},u))$.
 By taking the infimum of both sides of this inequality, we obtain
-    $\inf_{u\in U}N(f(x,u))\leq\inf_{u\in U}N(f(\overline{x}^{k},u))\leq\min_{j\in\mathcal{S}}N(f(\overline{x}^{k},\overline{u}^{j}))\leq\delta,$ (24)
+    (B.1) $\inf_{u\in U}N(f(x,u))\leq\inf_{u\in U}N(f(\overline{x}^{k},u))\leq\min_{j\in\mathcal{S}}N(f(\overline{x}^{k},\overline{u}^{j}))\leq\delta,$
     where the second inequality holds using the fact that $\{[\underline{u}^{i},\overline{u}^{i}]\}_{i\in\mathcal{S}}$ is a hyper-rectangular cover for $U$.
-The inequality (24) implies that $N$ satisfies condition (5).
-Therefore, the monotone neural network $N:\mathcal{X}\to$ satisfies conditions (2), (3), and (5). The result then follows using Theorem (2). This completes the proof of part 2.
-Appendix F Algorithm for Safety Verification
-In this appendix, we present an algorithm, analogous to Algorithm (1), for constructing monotone neural robust barrier certificates.
-Algorithm 2 Monotone Neural Robust Barrier Certificate
-Input: Simulator access to $f$, sets $\mathcal{X}_{0}$, $\mathcal{X}_{u}$, $\mathcal{X}=[\underline{x}_{\min},\overline{x}_{\max}]$, $e$ total number of epochs, $\hat{e}$ number of epochs before sampling, $m$ number of samples;
-Output: Neural network $B$ (barrier certificate);
-1:Over-approximate $\mathcal{X}_{0}$, $\mathcal{X}_{u}$, and $\mathcal{X}$ using unions of hyper-rectangles as in (6)
-2:Construct maximal initial points $\{\overline{x}^{i}\}_{i\in\mathcal{I}}$ and minimal unsafe points $\{\underline{x}^{i}\}_{i\in\mathcal{U}}$
-3:Construct the hyper-rectangular cover $T=[\underline{x}_{u},\overline{x}_{\max}]$
-4:Construct $H$ as the set of maximal points of $\mathcal{X}\setminus T$
-5:Initialize monotone neural networks $B$ and set hyperparameters $\gamma<\delta=\eta$
-6:for $e$ number of epochs do
-7:  Minimize loss $\widehat{L}_{\mathrm{total}}=L_{1}+L_{2}+L_{4}$ as defined in (11)- (13), respectively.
-8:  if Conditions (7)–(9) not satisfied for $\hat{e}$ epochs then
-9:   $Q\leftarrow$ Sample $m$ points from $\{x\in\mathcal{X}\mid\gamma<B(x)<\eta\}$
-10:   Update $T\leftarrow T\cup\bigcup_{x\in Q}[x,\overline{x}_{\max}]$
-11:   Update $H$ by getting maximal points of $\mathcal{X}\setminus T$
-12:   Set $\delta\leftarrow\min_{x\in H}N(x)$
-13:  end if
-14:  if all constraints are satisfied then
-15:   return trained network $B$
-16:  end if
-17:end for
-18:return Inconclusive (training failed to converge)
-Appendix G Correctness and Sample Complexity of Algorithm 1
-In this appendix, we provide additional details and proofs for Algorithm (1) presented in Section (3.3).
-We begin by presenting a systematic approach for computing the maximal elements of the set $\mathcal{X}\setminus T$.
-G.1 Computing Maximal Points of $\mathcal{X}\setminus T$
-[…]
-$T=\bigcup_{x\in Q\cup\{\underline{x}_{u}\}}[x,\overline{x}_{\max}].$ (25)
-Then, we can characterize the set $\mathcal{X}\setminus T$ as follows:
-$\mathcal{X}\setminus T=\left\{y\in\mathcal{X}\,\middle|\,\forall x\in Q\cup\{\underline{x}_{u}\},\,\exists i\in\{1,\ldots,n\}\text{ such that }y_{i}<x_{i}\right\}.$ (26)
-That is, a point $y\in\mathcal{X}$ lies in the region $\mathcal{X}\setminus T$ if it is strictly smaller than every $x\in Q\cup\{\underline{x}_{u}\}$ in at least one coordinate.
-To conservatively cover $\mathcal{X}\setminus T$, we construct a set $\mathcal{H}$ of maximal points defining hyper-rectangular regions. This is done through a structured procedure as follows:
-Step 1: Incomparable base sampling.
-We begin by sampling a set $Q=\{y^{1},\ldots,y^{m}\}\subset\mathcal{X}\setminus T$ such that all points are pairwise incomparable.
-These serve as the base points for constructing $\mathcal{H}$.
-Step 2: Maximal combinations. For each nonempty subset $S\subseteq\{y^{1},\ldots,y^{m}\}$, we compute the element-wise maximum
-$x_{S}:=\max\{y\mid y\in S\},$
-and collect the resulting points into a set $\mathcal{H}_{1}$. These represent the upper envelope of the base points.
-Step 3: Anchor points. To ensure full coverage of the domain boundaries, we define a set of anchor points. For each coordinate $i\in\{1,\ldots,n\}$, define $z^{i}\in\mathcal{X}$ by
-$(z^{i})_{j}=\begin{cases}(\overline{x}_{\max})_{j}&j\neq i\\ \min_{y\in Q}y_{i}&j=i\end{cases}$
-We collect the resulting points in the set $\mathcal{H}_{2}$, for every $i\in\{1,\ldots,n\}$.
-Step 4: Dominance filtering. We define the final set $\mathcal{H}$ as
-$\mathcal{H}:=\text{clipmax}(\mathcal{H}_{1}\cup\mathcal{H}_{2}),$ (27)
-where $\text{clipmax}(\cdot)$ denotes the subset of points not dominated by any other in the union.
-That is, we remove all $x\in\mathcal{H}_{1}\cup\mathcal{H}_{2}$ such that there exists $x^{\prime}\neq x$ with $x^{\prime}\leq x$.
-We will need the following Lemma to show that $\mathcal{H}$ consists of maximal points of the set $\mathcal{X}\setminus T$.
-Lemma 4 (Complement of a hyper-rectangle).
-Let $a\in\mathcal{X}$ and $A=[a,\overline{x}_{\max}]$ be a hyper-rectangle. Then we have
-$\mathcal{X}\setminus A\subseteq\bigcup_{y\in\mathcal{G}}[\underline{x}_{\min},y],$
-where $\mathcal{G}=\{a\}\cup\bigcup_{i=1}^{n}\{v^{i}\}$, where for every $i\in\{1,\ldots,n\}$, the point $v^{i}$ is defined as follows:
-$v^{i}_{j}=\begin{cases}(\overline{x}_{\max})_{j}&j\neq i\\ a_{i}&j=i\end{cases}$
-Proof.
-Let $w\in\mathcal{X}\setminus A$.
-This means that $w\not\geq a$.
-As a result, there exists $i\in\{1,\ldots,n\}$ such that $w_{i}<a_{i}$.
-Now consider the point $v^{i}$, for every $j\neq i$, we have $v^{i}_{j}=(\overline{x}_{\max})_{j}\geq w_{j}$.
-Moreover, we have $v^{i}_{i}=a_{i}>w_{i}$.
-This means that $v^{i}>w$ and therefore $w\in[\underline{x}_{\min},v^{i}]$.
-∎
-Now we extend the previous result for two hyper-rectangles.
-Lemma 5 (Complement of union of two hyper-rectangles).
-Let $a,b\in\mathcal{X}$ and $A=[a,\overline{x}_{\max}],B=[b,\overline{x}_{\max}]$ be two hyper-rectangles. Then we have
-$\mathcal{X}\setminus(A\cup B)\subseteq\bigcup_{y\in\mathcal{G}}[\underline{x}_{\min},y],$
-where $\mathcal{G}=\{x_{S}\}\cup\bigcup_{i=1}^{n}\{v^{i}\}$, where $x_{S}$ is element-wise maximum of $a$ and $b$ and for every $i\in\{1,\ldots,n\}$, the point $v^{i}$ is defined as follows:
-$v^{i}_{j}=\begin{cases}(\overline{x}_{\max})_{j}&j\neq i\\ \min\{a_{i},b_{i}\}&j=i\end{cases}$
-Proof.
-We know that $\mathcal{X}\setminus(A\cup B)=(\mathcal{X}\setminus A)\bigcap(\mathcal{X}\setminus B)$. By Lemma 4, we have
-$\mathcal{X}\setminus A\subseteq\bigcup_{y\in\mathcal{G}}[\underline{x}_{\min},y],\qquad\qquad\mathcal{X}\setminus B\subseteq\bigcup_{\widehat{y}\in\mathcal{\widehat{G}}}[\underline{x}_{\min},\widehat{y}$],
-where $\mathcal{G}=\{a\}\cup\bigcup_{i=1}^{n}\{v^{i}\}$, where for every $i\in\{1,\ldots,n\}$, the point $v^{i}$ is defined as follows:
-$v^{i}_{j}=\begin{cases}(\overline{x}_{\max})_{j}&j\neq i\\ a_{i}&j=i\end{cases}$
-and $\mathcal{\widehat{G}}=\{b\}\cup\bigcup_{i=1}^{n}\{\widehat{v}^{i}\}$, where for every $i\in\{1,\ldots,n\}$, the point $\widehat{v}^{i}$ is defined as follows:
-$\widehat{v}^{i}_{j}=\begin{cases}(\overline{x}_{\max})_{j}&j\neq i\\ b_{i}&j=i\end{cases}$
-Therefore, we have
-[…]
-$=\left([\underline{x}_{\min},a]\cup\bigcup_{i\in\{1,\ldots,n\}}[\underline{x}_{\min},v^{i}\}]\right)\bigcap\left([\underline{x}_{\min},b]\cup\bigcup_{i\in\{1,\ldots,n\}}[\underline{x}_{\min},\widehat{v}^{i}\}]\right)$
-$\subseteq\left([\underline{x}_{\min},x_{S}]\cup\bigcup_{i\in\{1,\ldots,n\}}[\underline{x}_{\min},v^{i}\}]\right)\bigcap\left([\underline{x}_{\min},x_{S}]\cup\bigcup_{i\in\{1,\ldots,n\}}[\underline{x}_{\min},\widehat{v}^{i}\}]\right)$
-$\subseteq[\underline{x}_{\min},x_{S}]\cup\left(\bigcup_{i\in\{1,\ldots,n\}}[\underline{x}_{\min},\min\{v^{i},\widehat{v}^{i}\}]\right),$
-where the first equality follows from Lemma 4, the third inequality follows from the interval inclusions $\underline{x}_{\min},a]\subseteq[\underline{x}_{\min},x_{S}]$ and $\underline{x}_{\min},b]\subseteq[\underline{x}_{\min},x_{S}]$, and the last inequality follows from the distributive property of unions and intersections.
-∎
-Lemma 5 can be extended to an arbitrary number of hyper-rectangles by applying mathematical induction.
-Theorem 6 (Maximal Points of $\mathcal{X}\setminus T$).
-Let $\mathcal{X}\subseteq\mathbb{R}^{n}$, and let $T=\bigcup_{x\in Q\cup\{\underline{x}_{u}\}}[x,\overline{x}_{\max}]$ be a finite union of upper hyper-rectangles.
-Suppose the set $\mathcal{H}$ is constructed through the procedure in Steps 1–4, as described in (27).
-Then
-$\mathcal{X}\setminus T\subseteq\bigcup_{y\in\mathcal{H}}[\underline{x}_{\min},y].$
-Proof.
-The results follows by applying Lemma 5 inductively. ∎
-Theorem 7 (Cardinality of $\mathcal{H}$).
-Let $\mathcal{X}\subseteq\mathbb{R}^{n}$, and let $T=\bigcup_{x\in Q\cup\{\underline{x}_{u}\}}[x,\overline{x}_{\max}]$ be a finite union of upper hyper-rectangles.
-Suppose the set $\mathcal{H}$ is constructed through the procedure in Steps 1–4, as described in (27).
-Then, the cardinality of the set $\mathcal{H}$ is $\mathcal{O}(n)$.
-Proof.
-By the procedure in Steps 1–4, we have $\mathcal{H}\subseteq\mathcal{H}_{1}\cup\mathcal{H}_{2}$.
-This implies that $|\mathcal{H}|\leq|\mathcal{H}_{1}|+|\mathcal{H}_{2}|$.
-Based on Step 2, we have $|\mathcal{H}_{1}|\leq 2^{m}-1$.
-Moreover, in Step 3, we construct $n$ anchor points.
-This means that $|\mathcal{H}_{2}|=n$.
-As a result, the cardinality of $\mathcal{H}$ is at most $2^{m}+n-1$ and therefore $\mathcal{O}(n)$.
-∎
-G.2 Proof of Theorem 5
-Proof.
-We divide the proof into two parts: (i) proof of correctness, and (ii) proof of linear sample complexity.
-  1. 1.
-    (Correctness).
-Suppose that Algorithm 1 converges and returns a monotone neural control barrier certificate $N$ and a feedback controller $K$.
-By the termination condition of Algorithm 1, $N$ must satisfy conditions (7)–(8).
-We now show that $N$ and $K$ also satisfy condition (9).
-Since the set $H$ consists of the maximal elements of $\mathcal{X}\setminus T$, and $\delta=\min_{x\in H}N(x)$, it follows that for every $x\in T$, we have $N(x)\geq\delta$.
-Therefore, condition (9) holds trivially on the region $T$ for the monotone neural control barrier certificate $N$ and the feedback controller $K$.
-Furthermore, the algorithm ensures that condition (9) also holds on $\mathcal{X}\setminus T$.
-Thus, $N$ and $K$ satisfy condition (9) over the entire state space $\mathcal{X}$.
-The result then follows directly from Theorem 4.
-  2. 2.
-    (Linear sample complexity).
-The sample complexity of Algorithm (1) is exactly the number of elements in the family of hyper-rectangles $\{[\underline{x}_{\min},\overline{x}^{i}]_{i\in\mathcal{P}}\}$ where $\overline{x}^{i}$ are the maximal points of the set $\mathcal{X}\setminus T$.
-By Theorem 7, the cardinality of the maximal points of $\mathcal{X}\setminus T$ is $|\mathcal{H}|$ and it is of $\mathcal{O}(n)$.
-This completes the proof of linear sample complexity.
-Appendix H Urban Traffic Control
-We consider an urban traffic network described by a family of links (roads) categorized according to their actuation type: $\mathcal{L}_{r}$ (roads with traffic light control), $\mathcal{L}_{o}$ (on-ramps controlled by ramp meters), and $\mathcal{L}_{f}$ (freeway segments without direct control).
-Uncontrolled roads are also treated like freeways, so the set of all links (roads) in this network is given by $\mathcal{L}=\mathcal{L}_{r}\cup\mathcal{L}_{o}\cup\mathcal{L}_{f}$.
-The number of vehicles on link $l$ at time $t$ is $x_{[l],t}\in[0,c_{l}]$, where $c_{l}$ is the capacity of link $l$.
-We assume that the traffic flow follows fluid-like behavior.
-Here, the state space is $\mathcal{X}:=\Pi_{l\in\mathcal{L}}[0,c_{l}]$.
-The outflow from link $l$ at time $t$ is:
-$q_{[l],t}=\min\Big{\{}x_{[l],t},\;\bar{q}_{l},\;\min_{l^{\prime}:\tau(l^{\prime})=\eta(l)}\frac{\alpha_{l:l^{\prime}}}{\beta_{l:l^{\prime}}}(c_{l^{\prime}}-x_{[l^{\prime}],t})\Big{\}},$ (28)
-where $\bar{q}_{l}$ is max outflow from $l$, $\alpha_{l:l^{\prime}}$ is capacity ratio of link $l^{\prime}$ relative to $l$, and $\beta_{l:l^{\prime}}$ is turning ratio from $l$ to $l^{\prime}$.
-Figure 2 presents a schematic diagram of the link outflow $q[l]$ as a function of the state variables $x[l]$ and $x[l^{\prime}]$, commonly referred to as the fundamental diagram in traffic flow literature [23].
-The flow $q[l]$ decreases when one or more downstream links lack sufficient capacity to accommodate incoming vehicles.
-In such cases—when the last term in (28) determines the minimum—we say the traffic is congested.
-Otherwise, if the upstream demand governs the flow, the traffic is considered to be in the free-flow regime.
-We define the congestion-free set $\Pi$ as:
-$\Pi:=\left\{x\in\mathcal{X}\mid\min\{x_{[l]},\bar{q}_{l}\}\leq\frac{\alpha_{l:l^{\prime}}}{\beta_{l:l^{\prime}}}(c_{l^{\prime}}-x_{[l^{\prime}]})\;\forall l,l^{\prime}:\tau(l^{\prime})=\eta(l)\right\}.$ (29)
-The actual allowable outflow from link $l$ depends on its actuation:
-$\tilde{q}_{[l],t}=\begin{cases}s_{[l],t}\cdot q_{[l],t}&\text{if }l\in\mathcal{L}_{r},\\ \min\{q_{[l],t},r_{[l],t}\}&\text{if }l\in\mathcal{L}_{o},\\ q_{[l],t}&\text{if }l\in\mathcal{L}_{f}.\end{cases}$ (30)
-Here, $s_{[l],t}\in\{0,1\}$ indicates the traffic light status (1 for green, 0 for red), and $r_{[l],t}\in\mathbb{R}_{+}$ is the ramp metering rate, both of which are control input to the system.
-Thus, $U=\mathbb{R}_{>0}^{4}\times\{0,1\}^{14}$.
-To prevent conflicting greens, for any pair $l,l^{\prime}$ with perpendicular intersections, we require $s_{[l],t}+s_{[l^{\prime}],t}\leq 1$.
-The state transition map $f$ for this urban traffic network is defined by:
-$x_{[l],t+1}=x_{[l],t}-\tilde{q}_{[l],t}+\sum_{l^{\prime}:\tau(l)=\eta(l^{\prime})}\beta_{l^{\prime}:l}\tilde{q}_{[l^{\prime}],t}:=f_{[l]}(x_{t},s_{t},r_{t}),\qquad\mbox{ for all }l\in\mathcal{L}.$ (31)
-We define the set of unsafe states as $\mathcal{X}_{u}=\mathcal{X}\setminus\Pi$; that is, a state is considered unsafe if at least one link in the network is congested.
-Some trajectories of the system are depicted in Figure 3.
-We refer the reader to [42] for more details on model of this system, its topology, and proof of monotonicty.
-Parameters of the model are found in Table 3
-$x_{[l]},x_{[l^{\prime}]}$$q_{[l]}$$\frac{\alpha_{l:l^{\prime}}}{\beta_{l:l^{\prime}}}(c_{l^{\prime}}-x_{[l^{\prime}]})$free flowcongested flow
-Figure 2: Fundamental diagram.
-The outflow from link $l$ decreases when the number of vehicles on its immediate downstream link $l^{\prime}$ approaches the downstream capacity.
-This reduction in flow, caused by limited downstream supply, characterizes the onset of congestion due to blocking behavior.
-Figure 3: Trajectories of the system. Dotted lines indicate entering congestion status.
-Table 3: Parameters of the urban traffic network
-links parameters
-1–26, 49–53 $\bar{q}=15$, $c=40$
-27–34 $\bar{q}=15$, $c=30$
-35–48 $\bar{q}=40$, $c=60$
-Turning ratios value
-$\beta_{2:50},\beta_{4,53},\beta_{8:51},\beta_{12:7},\beta_{13:28},\beta_{15:30},\beta_{16:28}$ 0.2
-$\beta_{21:32},\beta_{24:32},\beta_{26:2},\beta_{36:31},\beta_{36:33},\beta_{39:27},\beta_{43:29}$
-$\beta_{5:12},\beta_{6:13},\beta_{6:18},\beta_{10:21},\beta_{10:26}$ 0.3
-$\beta_{1:20},\beta_{6:7}$ 0.4
-$\beta_{1:2},\beta_{11:12},\beta_{14:30},\beta_{17:7},\beta_{17:18},\beta_{19:2},\beta_{19:20}$
-$\beta_{31:22},\beta_{31:25},\beta_{33:24},\beta_{49:3},\beta_{49:50},\beta_{51:4},\beta_{52:5}$ 0.5
-$\beta_{22:34},\beta_{23:24},\beta_{23:34},\beta_{27:14},\beta_{27:17},\beta_{29:16},\beta_{52:53}$
-$\beta_{2:3},\beta_{3:4},\beta_{4:5},\beta_{8:9},\beta_{12:13},\beta_{13:14},\beta_{15:16},\beta_{16:17}$
-$\beta_{20:21},\beta_{21:22},\beta_{24:25},\beta_{25:26},\beta_{36:37},\beta_{39:40}$ 0.8
-$\beta_{43:44},\beta_{46:47}$
-Capacity ratios value
-$\alpha_{19:2},\alpha_{26:2},\alpha_{17:7},\alpha_{12:7},\alpha_{13:28},\alpha_{16:28}$ $0.5$
-$\alpha_{14:30},\alpha_{15:30},\alpha_{21:32},\alpha_{24:32},\alpha_{22:34},\alpha_{23:34}$
-Appendix I Freeway Traffic Verification
-For the freeway traffic network $\mathfrak{S}=(\mathcal{X},\mathcal{X}_{0},U,f)$ shown in Figure (4), the state space is given by $\mathcal{X}=[0,50]^{n}$, and the state transition map $f:\mathcal{X}\times U\to\mathcal{X}$ is defined by:
-$f_{1}(q)$ $=q_{1}+\tau\left(F_{0\rightarrow 1}-F_{1\rightarrow 2}(q)\right),$
-$f_{i}(q)$ $=q_{i}+\tau\left(F_{i-1\rightarrow i}(q)-F_{i\rightarrow i+1}(q)\right),\quad i\in\{2,\ldots,n-1\},$
-$f_{n}(q)$ $=q_{n}+\tau\left(F_{n-1\rightarrow n}(q)-F_{n\rightarrow 0}(q)\right),$ (32)
-where the input is the on-ramp inflow $F_{0\rightarrow 1}\in U=[0,9]$, the parameter $\tau=0.1$ is sampling time, and $q=\begin{bmatrix}q_{1}&\ldots&q_{n}\end{bmatrix}^{\top}$ with $q_{i}\geq 0$ representing the vehicle density in road segment $i$.
-The flow from road segment $i$ to $j$, denoted $F_{i\rightarrow j}(q)$, is defined as:
-$F_{i\rightarrow j}(q)$ $=\min\left\{\varphi_{i}(q_{i}),\sigma_{j}(q_{j})\right\},$ (33)
-$F_{n\rightarrow 0}(q)$ $=\varphi_{n}(q_{n}),$ (34)
-where $\varphi_{i}(q_{i})$ models the increasing outflow demand of segment $i$, and $\sigma_{j}(q_{j})$ models the decreasing inflow supply of segment $j$.
-The demand and supply functions are given by:
-$\varphi_{i}(q_{i})$ $=15\left(1-e^{-0.038q_{i}}\right),\quad\sigma_{i}(q_{i})=\mathrm{ReLU}\left(15\left(1-\left(\frac{q_{i}}{100}\right)^{4}\right)\right),\quad\forall i\neq 2,$
-$\varphi_{2}(q_{2})$ $=10\left(1-e^{-0.138q_{2}}\right),\quad\sigma_{2}(q_{2})=\mathrm{ReLU}\left(10\left(1-\left(\frac{q_{2}}{50}\right)^{4}\right)\right).$ (35)
-The initial set for this system is defined as $\mathcal{X}_{0}=[0,5]\times[0,18]\times[0,5]^{n-2}$.
-The unsafe set is $\mathcal{X}_{u}=[30,50]^{n}$, which indicates congestion in all roads.
-We refer the reader to [9] for more details on this system and proof of monotonicity.
-For $n=1000$, our Algorithm (F) find a monotone neural robust barrier certificate for $\mathfrak{S}$ in 135s (averaged over 30 runs).
-Figure 4: A line digraph representing an $n$-dim traffic network.
+The inequality (B.1) implies that $N$ satisfies condition (2.5).
+Therefore, the monotone neural network $N:\mathcal{X}\to\mathbb{R}$ satisfies conditions (2.2), (2.3), and (2.5). The result then follows using Theorem (2.5). This completes the proof of part 2.